### v2

# Computation of Smyth and Hoare Power Constructions in Well-filtered Dcpos

Abstract
Prior work [11] established commutativity result for the Hoare power construction and a modified version of the Smyth power construction on all $\mathcal{U}_{S}$-admitting dcpos, where the $\mathcal{U}_{S}$-admissibility determined by the strongly compact subsets instead of compact subsets is weaker than the well-filteredness. In this paper, we introduce the Hoare power construction $\mathcal{H}$ and the Smyth power construction $\mathcal{Q}$ on the category $\mathbf{WF}$ of well-filtered dcpos with Scott continuous maps. Actually, the fubctors $\mathcal{H}$ and $\mathcal{Q}$ can be extended to monads. We prove that $\mathcal{H}$ and $\mathcal{Q}$ commute, that is, $\mathcal{HQ}(L)$ is isomorphic to $\mathcal{QH}(L)$ for a well-filtered dcpo $L$, if and only if $L$ satisfies the property $(\mathrm{KC})$ and the Scott topology coincides with the upper Vietoris topology on $\mathcal{Q}(L)$. We also investigate the Eilenberg-Moore category of the monad composed by $\mathcal{H}$ and $\mathcal{Q}$ on $\mathbf{WF}$ and characterize it to be a subcategory of the category $\mathbf{Frm}$, which is composed of all frames and all frame homomorphisms.
keywords:
Well-filtered dcpos , Power constructions , Monad , $\mathrm{Eilenbleg}$-$\mathrm{Moore}$ category , Commutativity
MSC:
18C15, 18C20
t1t1footnotetext: This work is supported by the National Natural Science Foundation of China (No.12231007)
1 Introduction
Non-deterministic choices are important semantic concepts that offer new insights in designing more powerful programming languages. They were originally modeled by the power constructions on the category of domains. The most common of these are the Hoare and the Smyth powerdomains [21], which can respectively serve as the denotational semantics for angelic and demoniac non-deterministic choices. Notably, the Hoare and the Smyth powerdomains also give rise to monads. In 1991, Eugenio Moggi showed that, within the category of sets, the collection of all finite sets could serve as denotations for nondeterministic programming languages [18]. He further proved that this structure forms a monad on the category of sets, thereby linking power constructions with monads. Over time, monads have evolved into a fundamental tool for constructing denotational semantic models in functional programming languages.
In 1969, Jon Beck proposed the notion of the distributive law between two monads defined on the same category [2]. Such a distributivity ensures that the composition of these two monads remains a monad, providing a new approach to establishing denotational semantic models in functional programming languages. Moreover, the two composite monads obtained under different composition orders may differ; if they coincide, i.e., the monad functors induce isomorphic mappings on objects, then the monads are said to obey a commutativity law. Such commutativity ensures that the order of program execution becomes immaterial in computation. This practical significance raises two fundamental problems: (1) Across which categories does a distributive law exist between the two denotational semantic models for nondeterministic programming languages-the Hoare power monad and the Smyth power monad? (2) Within a specified category, under what conditions are these two monads commutative?
In domain theory, numerous scholars have conducted in-depth investigations into the above two problems. In 1990, Kevin Flannery and Johannes Martin employed information systems to demonstrate the commutativity between the Hoare power monad and the Smyth power monad in the category of bounded complete algebraic domains [6]. Subsequently, Reinhold Heckmann used an entirely distinct method to prove that for any full subcategory of the category of dcpos, the commutativity of Hoare and Smyth power monads is guaranteed when the two monads can be characterized as the free inflationary semilattice and the free deflationary semilattice, respectively [12]. Within the category of domains, the Hoare powerdomain induces the free upper semilattices, while the Smyth powerdomain yields the free meet-continuous semilattices [9]. Here, the upper semilattices are actually inflationary semilattices, and the meet-continuous semilattices are deflationary semilattices. Thus the Hoare power monad and the Smyth power monad commute in the category of domains.
However, the fact remains: not all full subcategories of the category of dcpos admit power monads that are free semilattice constructions. This means that the results in [12] cannot be applied to verify the commutativity of the Smyth and Hoare power monads. A canonical example is the category of well-filtered dcpos, in which the Smyth power monad does not admit a free deflationary semilattice construction. Given that well-filteredness is a fundamentally important property in order theory, many scholars have conducted investigations into the commutativity between Hoare and Smyth power monads in the category of well-filtered dcpos. It is essential to note that Heckmann in 1991 investigated the commutativity between the Hoare power construction and a modified Smyth power construction on the category of $U_{S}$-admitting dcpos [11]. In this work, the extended Smyth power construction is determined by all strongly compact subsets, not the traditional Smyth power construction based on compact subsets. Thus, the $\mathcal{U}_{S}$-admissibility is strictly weaker than well-filteredness. Furthermore, Heckmann established only a sufficient (but not necessary) condition for the commutativity.
In this paper, we focus on the category of well-filtered dcpos and develop new techniques to investigate the commutativity between the associated Hoare and Smyth power monads. In Section 3, we define the Hoare and Smyth constructions for a well-filtered dcpo $L$ as follows: The Hoare construction $\mathcal{H}$ is the set of all Scott closed subsets of $L$, ordered by inclusion; the Smyth construction $\mathcal{Q}$ is the set of all nonempty Scott compact saturated subsets of $L$, ordered by reverse inclusion order. We explicitly show that $\mathcal{Q}$ does not yield a free construction for the category $\mathbf{WF}$ of well-filtered dcpos with Scott continuous maps, which confirms the significance and necessity of our research. In Section 4, we introduce a property called $(\mathrm{KC})$, which is equal to the consonance in topological spaces but weaker than that in terms of posets. Based on this we proved that $\mathcal{H}$ and $\mathcal{Q}$ commute in the sense of that $\mathcal{HQ}(L)$ is isomorphic to $\mathcal{QH}(L)$ under a certain function, if and only if $L$ satisfies the property $(\mathrm{KC})$ and the Scott topology coincides with the upper Vietoris topology on $\mathcal{Q}(L)$. Notably, unlike Heckmann’s modified Smyth power construction $\mathcal{U}_{S}$, our $\mathcal{Q}$ induces the Smyth power monad on $\mathbf{WF}$. This relies crucially on Xu’s result [24] that the Smyth power space endowed with the Scott topology of any well-filtered space remains well-filtered. Furthermore, since every complete lattice with the Scott topology is well-filtered, we conclude that $\mathcal{H}$ also forms a monad on $\mathbf{WF}$. Thus in Section 5, we analyze the properties of the composite monad. The morphisms $\phi$ and $\psi$ defined in Definition 4.1 have also shown to be distributive laws, ensuring that the composition of $\mathcal{H}$ and $\mathcal{Q}$, regardless of the order, is still a monad on $\mathbf{WF}$. Finally, motivated by growing interest in Eilenberg-Moore algebras, we characterize the Eilenberg-Moore category of $\mathcal{QH}$ (when isomorphic to $\mathcal{HQ}$ under $\phi$ and $\psi$) as a subcategory of Frm-the category of frames and frame homomorphisms.
2 Preliminaries
In this section, we will give a brief introduction to some basic concepts and notions to be used in this paper. One can refer to [1, 9, 10] for more details.
Given a poset $P$, $A\subseteq P$ is an upper set if $A={\uparrow}A$, where ${\uparrow}A=\{x\in P:x\geq a\ \mathrm{for\ some}\ a\in A\}$. The lower set is defined dually. The subset $D$ of $P$ is directed if every finite subset of $D$ has an upper bound in $D$. A subset $U$ of $P$ is called Scott open iff it is an upper set and for each directed subset $D$ whose supremum exists, written as $\sup D$, $\sup D\in U$ implies $D\cap U\neq\emptyset$. The complement of a Scott open set is called Scott closed. The collection of all Scott open subsets, denoted by $\sigma(P)$, form a topology called Scott topology and in general, we write $\Sigma P=(P,\sigma(P))$. Besides, we use $\Gamma(P)$ to denote the set of all Scott closed subsets of $P$. A continuous map between two posets equipped with Scott topology is called Scott-continuous.
Let $X$ be a topological space. When $X$ is a $T_{0}$ space, the specialization preorder $\leq$ defined by $x\leq y$ iff $x\in cl(\{y\})$ is a partial order and is called specialization order. A subset of $X$ is saturated if it is the intersection of open sets, or equivalently, it is an upper set in the order of specialization. $X$ is called well-filtered if for each filter basis $\mathcal{C}$ of compact saturated subsets and any open subset $U$ of $X$ with $\bigcap\mathcal{C}\subseteq U$, $K\subseteq U$ for some $K\in\mathcal{C}$. We usually say a dcpo $P$ well-filtered if it equipped with the Scott topology is a well-filtered space.
In this paper, we denote the set of all nonempty compact saturated subsets and that of all closed subsets of a space $X$ by $\mathcal{Q}(X)$ and $\Gamma(X)$, respectively. The topology on $\mathcal{Q}(X)$ generated by $\{\square U:U\in\mathcal{O}(X)\}$ as a base, where $\Box U=\{K\in Q(X):K\subseteq U\}$, is called the upper Vietoris topology and denoted by $\upsilon(Q(X))$. The upper topology on $\Gamma(X)$ coincides with the lower Vietoris topology, which is written as $\nu(\Gamma(X))$ and generated by $\{\lozenge U:U\in\mathcal{O}(X)\}$ as a subbase, where $\lozenge U=\{A\in\Gamma(X):A\cap U\neq\emptyset\}$.
Definition 2.1.
([1]) A monad $\mathcal{T}$ in a category $\mathbf{C}$ is a triple consists of an endofunctor $\mathcal{T}:\mathbf{C}\rightarrow\mathbf{C}$ and two natural transformations: the unit $\eta:Id_{\mathbf{C}}\rightarrow\mathcal{T}$ and the multiplication $\mu:\mathcal{T}^{2}\rightarrow\mathcal{T}$ such that the following diagrams commute.
$\mathcal{T}$$\mathcal{T}^{2}$$\mathcal{T}$$\mathcal{T}$$\eta_{\mathcal{T}}$$\mathcal{T}\mu$$\mu$==$\mathcal{T}^{3}$$\mathcal{T}^{2}$$\mathcal{T}^{2}$$\mathcal{T}$$\mathcal{T}\mu$$\mu\mathcal{T}$$\mu$$\mu$
Definition 2.2.
([1]) Let $\mathcal{T}=(\mathcal{T},\eta,\mu)$ be a monad on the category $\mathbf{C}$.
  1. 1.
    A pair $(C,\alpha)$ is a $\mathcal{T}$-algebra (or Eilenberg-Moore algebra of $\mathcal{T}$) if $\alpha:\mathcal{T}C\rightarrow C$ is a morphism in $\mathbf{C}$ that satisfies:
    ($\mathbf{Associativity}$) $\alpha\circ\mu_{C}=\alpha\circ\mathcal{T}\alpha$,  ($\mathbf{Unit\;law}$) $\alpha\circ\eta_{C}=id_{C}$.
    In this case, $\alpha$ is called a structure map.
  2. 2.
    Let $(A,\alpha_{A})$ and $(B,\alpha_{B})$ be $\mathcal{T}$-algebras. A $\mathcal{T}$-algebra homomorphism from $(A,\alpha_{A})$ to $(B,\alpha_{B})$ is an arrow $h:A\rightarrow B$ satisfying:
    $(\mathbf{Homomorphism\;law})\ h\circ\alpha_{A}=\alpha_{B}\circ\mathcal{T}h$.
The category composed of $\mathcal{T}$-algebras and their homomorphisms is called Eilenberg-Moore category of $\mathcal{T}$ over $\mathbf{C}$.
3 The Hoare and Smyth power constructions
3.1 The Hoare power construction
Definition 3.1.
The Hoare power construction $\mathcal{H}(L)$ for a well-filtered dcpo $L$ is the set of all Scott closed subsets of $L$ ordered by inclusion.
We now focus on constructing mutually adjoint functors, through which the composite functor induces the monad $\mathcal{H}$ on $\mathbf{WF}$. Clearly, for each well-filtered dcpo $L$, its Hoare power construction is a complete lattice. Let $\mathbf{CL}$ denote the category of complete lattices with morphisms preserving arbitrary suprema.
Consider the morphisms in $\mathbf{WF}$. For any Scott continuous mapping $f:L\rightarrow M$ between well-filtered dcpos $L$ and $M$, we define $\mathcal{H}(f):\mathcal{H}(L)\rightarrow\mathcal{H}(M)$ as
$\mathcal{H}(f)(A)=cl(f(A))$, for any $A\in\mathcal{H}(L)$.
Obviously, for any collection $\{A_{i}:i\in I\}\subseteq\mathcal{H}(L)$,
$\mathcal{H}(f)(\sup_{i\in I}A_{i})=\mathcal{H}(f)(cl(\bigcup_{i\in I}A_{i}))=cl(f(\bigcup_{i\in I}A_{i}))=cl(\bigcup_{i\in I}f(A_{i}))=\sup_{i\in I}\mathcal{H}(f)(A_{i})$.
This means that $\mathcal{H}(f)$ preserves arbitrary suprema. Additionally, one can verify that $\mathcal{H}$ preserves the identities and the compositions of any two morphisms. Along with the above, we can obtain the following result.
Lemma 3.2.
$\mathcal{H}$ is a functor from $\mathbf{WF}$ to $\mathbf{CL}$.
It follows [23, Corollary 3.2] that each complete lattice endowed with the Scott topology is well-filtered. So there is a forgetful functor from $\mathbf{CL}$ to $\mathbf{WF}$.
Proposition 3.3.
The functor $\mathcal{H}:\mathbf{WF}\rightarrow\mathbf{CL}$ is left adjoint to the forgetful functor $U:\mathbf{CL}\rightarrow\mathbf{WF}$.
Proof.
We need to show that for each well-filtered dcpo $L$, its Hoare power construction $\mathcal{H}(L)$ has the universal property, that is, for any complete lattice $M$ and any Scott continuous map $f:L\rightarrow U(M)$, there is a unique morphism $\hat{f}:\mathcal{H}(L)\rightarrow M$ in $\mathbf{CL}$ such that $f=U(\hat{f})\circ\eta_{L}$, where $\eta_{L}:L\rightarrow U(\mathcal{H}(L))$ is an embedding map defined as
$\eta_{L}(x)={\downarrow}x$, for any $x\in L$.
Define $\hat{f}:\mathcal{H}(L)\rightarrow M$ as
$\hat{f}(A)=\sup f(A)$, for any $A\in\mathcal{H}(L)$,
whose rationality is guaranteed by the completeness of $M$. One can easily verify that $\hat{f}$ preserves arbitrary sups and $U(\hat{f})\circ\eta_{L}(x)=\sup f({\downarrow}x)=f(x)$ for all $x\in L$. From the fact that $A=\bigcup_{a\in A}{\downarrow}a=\sup_{\mathcal{H}(L)}\{{\downarrow}a:a\in A\}$ and the requirement that $\hat{f}$ should be in $\mathbf{CL}$, we know the definition of $\hat{f}$ can only be given in this way. ∎
Since each pair of adjunction determines a monad, one can refer to [1, Proposition 4.2.1], we deduce that the triple $(U\circ\mathcal{H},\eta,U\varepsilon_{\mathcal{H}})$, where $\eta$ and $\varepsilon$ are the unit and counit respectively, turns into a monad on $\mathbf{WF}$. For ease of expressions, we will abbreviate $U\circ\mathcal{H}$ as $\mathcal{H}$. Meanwhile, when $U\varepsilon_{\mathcal{H}}$ works on a well-filtered $\mathrm{dcpo}$ $L$, one can calculate $U\varepsilon_{\mathcal{H}(L)}$ as the natural transformation from $\mathcal{H}(\mathcal{H}(L))$ to $\mathcal{H}(L)$ that maps $\mathcal{A}$ to $\sup_{\mathcal{H}(L)}\mathcal{A}$.
Lemma 3.4.
Let $L$ be a dcpo and $\mathcal{A}$ a Scott closed subset of $\mathcal{H}(L)$. Then $\bigcup\mathcal{A}$ is Scott closed.
Proof.
It is similar to that of [14, Proposition 2.2]. ∎
It follows from this lemma that $\sup_{\mathcal{H}(L)}\mathcal{A}=\bigcup\mathcal{A}$. Now we conclude that
Theorem 3.5.
The endofunctor $\mathcal{H}$ together with the unit $\eta:Id\rightarrow U\circ\mathcal{H}$ and the multiplication $\mu:\mathcal{H}^{2}\rightarrow\mathcal{H}$ forms a monad on $\mathbf{WF}$, where $\eta$ and $\mu$ are defined concretely for a well-filtered dcpo $L$ as follows:
for any $x\in L,\eta_{L}(x)={\downarrow}x$,
and
for any $\mathcal{A}\in\mathcal{H}(\mathcal{H}(L)),\mu_{L}(\mathcal{A})=\bigcup\mathcal{A}$.
Combining the fact that $\mathbf{CL}$ is the category of unital inflationary semilattices for $\mathbf{WF}$ with the fact that the Hoare power well-filtered dcpo has the universal property, we have that
Corollary 3.6.
$\mathcal{H}$ gives the free unital inflationary semilattice for $\mathbf{WF}$.
Lemma 3.7.
Let $L,M$ be well-filtered dcpos. Then
  1. 1.
    $(L,\alpha)$ is an $\mathcal{H}$-algebra if and only if $L$ is a complete lattice.
  2. 2.
    $f:(L,\alpha)\rightarrow(M,\beta)$ is an $\mathcal{H}$-algebra homomorphism if and only if $f$ is a complete lattice homomorphism.
Proof.
Based on their definitions, it is easy to verify them. ∎
Corollary 3.8.
The Eilenberg-Moore category of $\mathcal{H}$ over $\mathbf{WF}$ is just $\mathbf{CL}$.
3.2 The Smyth power construction
Xu et al. proved that for a well-filtered space $X$, $\Sigma\mathcal{Q}(X)$ is still well-filtered [24], which indicates that for each well-filtered dcpo $L$, $\mathcal{Q}(L)$ with the Scott topology is well-filtered, that is, $\mathcal{Q}(L)$ is an object in $\mathbf{WF}$. This fact makes us give a rational Smyth power construction.
Definition 3.9.
The Smyth power construction for a well-filtered dcpo is the set of all nonempty Scott compact saturated subsets ordered by the reverse inclusion.
We shall use a direct way being different from that of $\mathcal{H}$ to establish $\mathcal{Q}$ as a monad on $\mathbf{WF}$.
Lemma 3.10.
Let $f:L\rightarrow M$ be a Scott-continuous mapping between well-filtered dcpos $L$ and $M$. Then $\mathcal{Q}(f):\mathcal{Q}(L)\rightarrow\mathcal{Q}(M)$ defined as
$\mathcal{Q}(f)(K)={\uparrow}f(K)$, for any $K\in\mathcal{Q}(L)$
is Scott-continuous.
Proof.
Let $\{K_{i}:i\in I\}$ be a filtered family in $\mathcal{Q}(L)$ with $\sup_{i\in I}K_{i}$ existing, that is to say, $\bigcap_{i\in I}K_{i}\in\mathcal{Q}(L)$. Since $f$ is Scott-continuous, $\{{\uparrow}f(K_{i}):i\in I\}$ is a subset of $\mathcal{Q}(M)$ and filtered. By the well-filteredness of $M$, we know $\bigcap_{i\in I}{\uparrow}f(K_{i})\in\mathcal{Q}(M)$, which means $\sup_{i\in I}{\uparrow}f(K_{i})$ exists in $\mathcal{Q}(M)$. Now we just need to prove that $\mathcal{Q}(f)(\bigcap_{i\in I}K_{i})=\bigcap_{i\in I}{\uparrow}f(K_{i})$, more precisely, to show
$\bigcap_{i\in I}{\uparrow}f(K_{i})\subseteq\mathcal{Q}(f)(\bigcap_{i\in I}K_{i})={\uparrow}f(\bigcap_{i\in I}K_{i})$
holds since the other containment is clear. If there is an element $x\in\bigcap_{i\in I}{\uparrow}f(K_{i})$ but $x\notin{\uparrow}f(\bigcap_{i\in I}K_{i})$, then $\bigcap_{i\in I}K_{i}$ will be contained in $f^{-1}(M\setminus{\downarrow}x)\in\sigma(L)$. Due to the well-filteredness of $L$, there must be an $i_{0}\in I$ such that $K_{i_{0}}\subseteq f^{-1}(M\setminus{\downarrow}x)$. Thus $f(K_{i_{0}})\subseteq M\setminus{\downarrow}x$ and so $x\notin{\uparrow}f(K_{i_{0}})$, which contradicts the assumption that $x\in\bigcap_{i\in I}{\uparrow}f(K_{i})$. ∎
One can easily verify that $\mathcal{Q}$ preserves the identities and the compositions of any two morphisms. So, $\mathcal{Q}$ is an endofunctor on $\mathbf{WF}$.
Lemma 3.11.
Let $L$ be a well-filtered dcpo and $\mathcal{K}$ a Scott compact saturated subset of $\mathcal{Q}(L)$. Then $\bigcup\mathcal{K}$ is compact saturated in $\Sigma L$.
Proof.
The saturation is clear. We only consider the compactness. Let $\mathcal{U}$ be a directed subset of $\sigma(L)$ with $\bigcup\mathcal{K}\subseteq\bigcup\mathcal{U}$. Then for each $K\in\mathcal{K}$, its compactness yields the existence of $U_{K}\in\mathcal{U}$ such that $K\subseteq U_{K}$, that is, $K\in\Box U_{K}$. Since the well-filteredness of $L$ hints that each $\Box U$ is Scott open in $\mathcal{Q}(L)$ and $\mathcal{U}$ is directed, $\{\Box U:U\in\mathcal{U}\}$ forms a directed family of Scott open subsets of $\mathcal{Q}(L)$ containing $\mathcal{K}$. So $\mathcal{K}\subseteq\Box U_{0}$ for some $U_{0}\in\mathcal{U}$ by its compactness. Thus $\bigcup\mathcal{K}\subseteq U_{0}$. ∎
Theorem 3.12.
The endofunctor $\mathcal{Q}$ with the unit $\theta$ and the multiplication $\iota$ forms a monad on $\mathbf{WF}$, where $\theta$ and $\iota$ are defined concretely as follows when they work on a well-filtered dcpo $L$,
$\theta_{L}(x)={\uparrow}x$, for any $x\in L$,
and
$\iota_{L}(\mathcal{K})=\bigcup\mathcal{K}$, for any $\mathcal{K}\in\mathcal{Q}((\mathcal{Q}(L)))$.
Proof.
It is a routine to check that $\mathcal{Q}$ is a monad. ∎
It is noted that for a well-filtered dcpo $L$, arbitrary finite infs and directed sups in $\mathcal{Q}(L)$ are actually the finite unions and filtered intersections. Thus $\mathcal{Q}(L)$ is a meet continuous semilattice, in other words, $\mathcal{Q}(L)$ is a deflationary semilattice in $\mathbf{WF}$.
Lemma 3.13.
Every $\mathcal{Q}$-algebra $M$ is a deflationary semilattice in $\mathbf{WF}$.
Proof.
The definitions of $\mathcal{Q}$-algebras reveal that $M$ is a retract of $\mathcal{Q}(M)$. Then by [20, proposition 5.6], we know that $M$ is a deflationary semilattice in $\mathbf{WF}$.
Lemma 3.14.
Let $(M,\alpha)$ be a $\mathcal{Q}$-algebra. Then for each $K\in\mathcal{Q}(M)$, its infimum $\wedge K$ exists.
Proof.
It is easy to prove that $\alpha(K)=\wedge K$ by utilizing the condition $\alpha\circ\theta_{M}=id_{M}$. ∎
Lemma 3.15.
Each $\mathcal{Q}$-algebra homomorphism preserves finite infs.
Proof.
Let $h:(L,\alpha_{L})\rightarrow(M,\alpha_{M})$ be a $\mathcal{Q}$-algebra homomorphism. Then we have $h\circ\alpha_{L}(K)=\alpha_{M}\circ\mathcal{Q}(h)(K)$ for any $K\in\mathcal{Q}(L)$, that is, $h(\wedge K)=\wedge h(K)$ holds. Accordingly, one can deduce that $h$ preserves finite infs as every finitely generated subset is compact saturated. ∎
Proposition 3.16.
The Eilenberg-Moore category of $\mathcal{Q}$ over $\mathbf{WF}$ is a subcategory of the category of deflationary semilattices in $\mathbf{WF}$.
Unfortunately, we can not accurately characterize what the $\mathcal{Q}$-algebras are. What is more, $\mathcal{Q}$ can not even give the free deflationary semilattice for $\mathbf{WF}$. Once $\mathcal{Q}$ can, each free deflationary semilattice in $\mathbf{WF}$ together with a structure map will be a $\mathcal{Q}$-algebra, which is violated in the following counterexample.
Example 3.17.
Let $\mathcal{L}$ be the dcpo given in Section 7.2.4 by Schalk in [20], that is,
=$\mathcal{L}=\{A\subseteq[0,1]:A\neq\emptyset\ \mathrm{and}\ [0,1]\setminus A\ \mathrm{is\ countable}\}$.
It has been proved to be a deflationary semilattice in $\mathbf{DCPO}$. Besides, we can see that every nonempty finite subset of $\mathcal{L}$ has a sup, which is indeed the union of its elements. It then follows that $\mathcal{L}$ is an up-complete $\vee$-semilattice, this implies the well-filteredness of $\Sigma\mathcal{L}$ witnessed by [23, Proposition 3.1]. Thus $\mathcal{L}$ is a deflationary semilattice in $\mathbf{WF}$. However, there exists a compact saturated subset
$K=\{[0,1]\setminus\{x\}:x\in[0,1/2]\}$,
which has no infimum in $\mathcal{L}$.
Hence, by Lemma 3.14, $\mathcal{L}$ fails to be a $\mathcal{Q}$-algebra. So $\mathcal{Q}$ can not give the free deflationary semilattice for $\mathbf{WF}$.
4 The commutativity between $\mathcal{H}$ and $\mathcal{Q}$
Definition 4.1.
Let $L$ be a well-filtered dcpo. We define $\phi_{L}:\mathcal{H}(\mathcal{Q}(L))\rightarrow\mathcal{Q}(\mathcal{H}(L))$ and $\psi_{L}:\mathcal{Q}(\mathcal{H}(L))\rightarrow\mathcal{H}(\mathcal{Q}(L))$ as
$\phi_{L}(\mathcal{A})=\{A\in\mathcal{H}(L):\forall K\in\mathcal{A},A\cap K\neq\emptyset\}$, for any $\mathcal{A}\in\mathcal{H}(\mathcal{Q}(L))$,
$\psi_{L}(\mathcal{K})=\{K\in\mathcal{Q}(L):\forall A\in\mathcal{K},K\cap A\neq\emptyset\}$, for any $\mathcal{K}\in\mathcal{Q}(\mathcal{H}(L))$.
For convenience, we will omit the subscripts when the object acted on is clear.
Lemma 4.2.
Every complete lattice endowed with the Scott topology is coherent.
Proof.
In [23], Corollary 3.2 guarantees the well-filteredness of a complete lattice $L$. Since for any $x,y\in L$, ${\uparrow}x\cap{\uparrow}y={\uparrow}(x\vee y)$, by Corollary 3.2 in [15], we have that $L$ is coherent. ∎
Lemma 4.3.
The functions $\phi$ and $\psi$ are well-defined.
Proof.
To make it clear, let us define $\phi^{\prime}:\mathcal{Q}(L)\rightarrow\mathcal{Q}(\mathcal{H}(L))$ and $\psi^{\prime}:\mathcal{H}(L)\rightarrow\mathcal{H}(\mathcal{Q}(L))$ as follows:
for any $K\in\mathcal{Q}(L)$, $\phi^{\prime}(K)=\{A\in\mathcal{H}(L):A\cap K\neq\emptyset\}$,
for any $A\in\mathcal{H}(L)$, $\psi^{\prime}(A)=\{K\in\mathcal{Q}(L):K\cap A\neq\emptyset\}$.
Then $\phi(\mathcal{A})=\bigcap\{\phi^{\prime}(K):K\in\mathcal{A}\}$ and $\psi(\mathcal{K})=\bigcap\{\psi^{\prime}(A):A\in\mathcal{K}\}$. One can easily check that $\phi^{\prime}(K)={\uparrow}\eta_{L}(K)$ and $\psi^{\prime}(A)=cl(\theta_{L}(A))$, which implies that $\phi^{\prime}$ and $\psi^{\prime}$ are well-defined. We know the intersection of arbitrary closed subsets is still closed. So, $\psi$ is well-defined. Now we focus on $\phi$. Lemma 4.2 reveals that $\Sigma\mathcal{H}(L)$ is coherent. Combining with the well-filteredness of $\Sigma\mathcal{H}(L)$, it will be a fact that each intersection of nonempty collection of compact saturated subsets in $\mathcal{H}(L)$ is still compact saturated. Thus $\phi(\mathcal{A})\in\mathcal{Q}(\mathcal{H}(L))$, i.e., $\phi$ is well-defined.
Definition 4.4.
A poset $P$ ($T_{0}$ space $X$) is said to have property $(\mathrm{KC})$ if for every compact saturated subset $\mathcal{K}$ of $\Gamma(P)$ under the Scott topology ($\mathcal{K}$ of $\Gamma(X)$ with respect to the lower Vietoris topology), any $U\in\sigma(P)$ ($U\in\mathcal{O}(X)$) that meets all members of $\mathcal{K}$ contains a compact saturated subset $K$ that still meets all members of $\mathcal{K}$.
Recall that a space $X$ is consonant if and only if for any Scott open subset $\mathcal{U}$ in $\mathcal{O}(X)$ and any $U\in\mathcal{U}$, there exists a $K\in Q(X)$ such that $U\in\Phi(K)\subseteq\mathcal{U}$, where $\Phi(K)=\{V\in\mathcal{O}(X):K\subseteq V\}$. The concept was proposed initially by Dolecki et al. in [5] to answer some questions related to topologies on the hyperspace of closed subsets of a topological space. We say a poset consonant if it endowed with the Scott topology is consonant.
Lemma 4.5.
If a poset $P$ is consonant, then it satisfies the property $(\mathrm{KC})$.
Proof.
Let $\mathcal{K}$ be a compact saturated subset of $(\Gamma(P),\sigma(\Gamma(P)))$ and $U$ a Scott open subset of $P$ satisfying $U\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. Set $\mathcal{U}=\bigcap\{\vartriangle\!\!A:A\in\mathcal{K}\}$, where $\vartriangle\!\!A=\{V\in\sigma(P):V\cap A\neq\emptyset\}$. We claim that $\mathcal{U}\in\sigma(\sigma(P))$. Obviously, it is an upper set. Assume $\mathcal{V}$ is a directed subset of $\sigma(P)$ and $\sup\mathcal{V}=\bigcup\mathcal{V}\in\mathcal{U}$. Then $\bigcup\mathcal{V}\in\vartriangle\!\!A$ for all $A\in\mathcal{K}$, that is, $\bigcup\mathcal{V}\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. It means that for each $A\in\mathcal{K}$, there exists $V_{A}\in\mathcal{V}$ such that $V_{A}\cap A\neq\emptyset$, i.e., $A\in\lozenge V_{A}$, and thus $\mathcal{K}\subseteq\bigcup\{\lozenge V:V\in\mathcal{V}\}$. By compactness of $\mathcal{K}$ and directedness of $\{\lozenge V:V\in\mathcal{V}\}$, there must exist a $V_{0}\in\mathcal{V}$ such that $\mathcal{K}\subseteq\lozenge V_{0}$. It follows that $V_{0}\in\vartriangle\!\!A$ for each $A\in\mathcal{K}$, that is to say, $V_{0}\in\mathcal{U}$. Thus $\mathcal{U}$ is Scott open. Since $P$ is consonant and $U\in\mathcal{U}$, there exists a $K\in Q(P)$ satisfying $U\in\Phi(K)\subseteq\mathcal{U}$. We will complete the proof if we show that $K\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. Suppose $K\cap A_{0}=\emptyset$ for some $A_{0}\in\mathcal{K}$. Then $K\subseteq P\setminus A_{0}$, which implies $P\setminus A_{0}\in\Phi(K)\subseteq\mathcal{U}$; so $P\setminus A_{0}\in\vartriangle\!\!A_{0}$, a contradiction. ∎
Lemma 4.6.
A $T_{0}$ space $X$ is consonant if and only if it has the property $(\mathrm{KC})$.
Proof.
()($\Rightarrow$)( ⇒ ): Let $\mathcal{K}$ be a compact saturated subset of $(\Gamma(X),\nu(\Gamma(X)))$ and $U\in\mathcal{O}(X)$ satisfying $U\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. Set $\mathcal{U}=\bigcap\{\blacktriangle A:A\in\mathcal{K}\}$, where $\blacktriangle A=\{V\in\mathcal{O}(X):V\cap A\neq\emptyset\}$. Then similar to the proof of Lemma 4.5, one can easily verify that $\mathcal{U}\in\sigma(\mathcal{O}(X))$ and find some $K\in Q(X)$ qualified.
()($\Leftarrow$)( ⇐ ): Let $\mathcal{V}$ be a Scott open subset of $\mathcal{O}(X)$ and $V\in\mathcal{V}$. We set $\mathcal{K}=\bigcap\{\lozenge U:U\in\mathcal{V}\}$. Notably, $\mathcal{K}$ is indeed the image of $\mathcal{V}$ under the function $\psi_{X}:\sigma(\mathcal{O}(X))\rightarrow\mathcal{Q}(\Gamma(X))$ given in [4, Definition 6.7], where $\mathcal{Q}(\Gamma(X))$ is the set of all compact saturated subset of $\Gamma(X)$ with respect to the lower Vietoris topology. Thus $\mathcal{K}$ is a compact saturated subset of $(\Gamma(X),\nu(\Gamma(X)))$. As $V\in\mathcal{V}$ suggests that $\mathcal{K}\subseteq\lozenge V$, $X\setminus V\notin\mathcal{K}$, more precisely, each $A$ in $\mathcal{K}$ is not contained in $X\setminus V$, that is, $V\cap A\neq\emptyset$ for each $A\in\mathcal{K}$. Since $X$ has the property $(\mathrm{KC})$, we can find a $K\in Q(X)$ contained in $V$ that intersects each $A\in\mathcal{K}$. To make the proof complete, we only need to show that $\Phi(K)\subseteq\mathcal{V}$. For each $W\in\Phi(K)$, the fact $K\subseteq W$ yields that $W\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. So $\mathcal{K}\in\Box\lozenge W$, which is equivalent to saying $W\in\phi_{X}({\mathcal{K}})$, where $\phi_{X}:\mathrm{K}(\Gamma(X))\rightarrow\sigma(\mathcal{O}(X))$ is also given in [4, Definition 6.7]. Moreover, [4, Theorem 6.10] has shown that $\phi_{X}\circ\psi_{X}=id_{\sigma(\mathcal{O}(X))}$, which enables us to deduce that $W\in\phi_{X}({\mathcal{K}})=\phi_{X}(\psi_{X}(\mathcal{V}))=\mathcal{V}$. Hence, we have $V\in\Phi(K)\subseteq\mathcal{V}$, which indicates the consonance of $X$.
Now see [4, Theorem 6.10] again, it states that for a space $X$, $\sigma(\mathcal{O}(X))$ and $\mathcal{Q}(\Gamma(X))$ are isomorphic lattices. Recently, Chen et al. in [3] proved that if a poset $P$ under the Scott topology is core-compact, then $\sigma(\Gamma(P))=\nu(\Gamma(P))$. Thus for a poset $P$, $\sigma(\sigma(P))$ would be isomorphic to $Q(\Gamma(P))$ as long as $\Sigma P$ is core-compact. Combining all of these results and the above two lemmas, we obtain the following result.
Corollary 4.7.
Let $P$ be a poset. If $(P,\sigma(P))$ is core-compact and has the property $(\mathrm{KC})$, then $P$ is consonant.
Theorem 4.8.
Let $L$ be a well-filtered dcpo. Then $\mathcal{H}(\mathcal{Q}(L))$ is isomorphic to $\mathcal{Q}(\mathcal{H}(L))$ under the maps $\phi$ and $\psi$ if and only if the following conditions are satisfied.
  1. 1.
    $\sigma(Q(L))\subseteq\upsilon(Q(L));$ and
  2. 2.
    $L$ has the property $(\mathrm{KC})$.
Proof.
()($\Rightarrow$)( ⇒ ): Let $\mathcal{U}$ be a Scott open subset of $\mathcal{Q}(L)$ and $K$ an element in $\mathcal{U}$. Then $K$ does not belong to the Scott closed subset $\mathcal{Q}(L)\setminus\mathcal{U}$. Since $\psi$ is surjective, there exists a $\mathcal{K}\in\mathcal{Q}(\mathcal{H}(L))$ such that $\mathcal{Q}(L)\setminus\mathcal{U}=\psi(\mathcal{K})$, then $K\notin\psi(\mathcal{K})$. By the definition of $\psi$, it yields that $K\cap A=\emptyset$ for some $A\in\mathcal{K}$. Now we find a Scott open subset $V=L\setminus A$ containing $K$, that is, $K\in\Box V$. Thus we only need to show $\Box V\subseteq\mathcal{U}$. For any $M\in\Box V$, $M\cap A=\emptyset$, which implies that $M\notin\psi(\mathcal{K})$. Hence, $M\in\mathcal{U}$.
Let $\mathcal{K}$ be a Scott compact saturated subset of $\Gamma(L)$ and $U\in\sigma(L)$ satisfying $U\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. Then $L\setminus U\notin\mathcal{K}$; otherwise, by the saturation of $\mathcal{K}$, there must be an $A_{0}\in\mathcal{K}$ such that $A_{0}\subseteq L\setminus U$, a contradiction. The fact that $\phi$ is surjective suggests that there is an $\mathcal{A}\in\mathcal{H}(\mathcal{Q}(L))$ such that $\mathcal{K}=\phi(\mathcal{A})$. Thus $L\setminus U\notin\phi(\mathcal{A})$. This means $(L\setminus U)\cap K_{0}=\emptyset$ i.e., $K_{0}\subseteq U$ for some $K_{0}\in\mathcal{A}$. Now suppose that $K_{0}\cap A_{1}=\emptyset$ for some $A_{1}\in\mathcal{K}$. Then $A_{1}\notin\phi(\mathcal{A})=\mathcal{K}$, which is an contradiction obviously. Thus $K_{0}$ intersects each $A\in\mathcal{K}$ and $L$ has the property $(\mathrm{KC})$ as desired.
()($\Leftarrow$)( ⇐ ): Obviously, $\phi$ and $\psi$ are order-preserving. Thus this proof will be completed so long as we can check that $\psi\circ\phi$ and $\phi\circ\psi$ are identities, more precisely, the following two equations hold (for arbitrary $\mathcal{A}\in\mathcal{H}(\mathcal{Q}(L))$ and $\mathcal{K}\in\mathcal{Q}(\mathcal{H}(L))$):
$\psi\circ\phi(\mathcal{A})=\{K\in\mathcal{Q}(L):\forall A\in\phi(\mathcal{A}),K\cap A\neq\emptyset\}=\mathcal{A}$,
$\phi\circ\psi(\mathcal{K})=\{A\in\mathcal{H}(L):\forall K\in\psi(\mathcal{K}),A\cap K\neq\emptyset\}=\mathcal{K}$.
Note that the definitions of $\phi$ and $\psi$ determine that $\mathcal{A}\subseteq\psi\circ\phi(\mathcal{A})$ and $\mathcal{K}\subseteq\phi\circ\psi(\mathcal{K})$. Thus we just need to check the reverse.
As to the first one, suppose that there is a $K\in\mathcal{Q}(L)$ which belongs to $\psi\circ\phi(\mathcal{A})$ but not to $\mathcal{A}$. Then $K\in\mathcal{Q}(L)\setminus\mathcal{A}$, which is Scott open in $\mathcal{Q}(L)$ and further an open subset of $\mathcal{Q}(L)$ with respect to the upper Vietoris topology by $(\mathrm{i})$. Thus there exists a $U\in\sigma(L)$ such that $K\in\Box U\subseteq\mathcal{Q}(L)\setminus\mathcal{A}$, which implies $K\cap(L\setminus U)=\emptyset$. This will contradict the fact that $K\in\psi\circ\phi(\mathcal{A})$ if we could show $L\setminus U\in\phi(\mathcal{A})$. Assume $(L\setminus U)\cap K_{0}=\emptyset$ for some $K_{0}\in\mathcal{A}$. Then $K_{0}\in\Box U$, thus $K_{0}\notin\mathcal{A}$, a contradiction. So $L\setminus U\in\phi(\mathcal{A})$. Therefore, $\psi\circ\phi(\mathcal{A})\subseteq\mathcal{A}$.
Regarding the second, for the sake of a contradiction, we assume that there is an $A\in\phi\circ\psi(\mathcal{K})$ but $A\notin\mathcal{K}$. It suggests that each $C\in\mathcal{K}$ is not contained in $A$, that is to say, $(L\setminus A)\cap C\neq\emptyset$ for all $C\in\mathcal{K}$. Since $L$ has the property $(\mathrm{KC})$, we can find a $K_{0}\subseteq L\setminus A$ intersecting each $C$ in $\mathcal{K}$. Then $K_{0}\in\psi(\mathcal{K})$ and $K_{0}\cap A=\emptyset$, which contradicts that $A\in\phi\circ\psi(\mathcal{K})$. Thus $A\in\mathcal{K}$, which obviously leads to a contradiction. So our assumption is wrong and $\phi\circ\psi(\mathcal{K})\subseteq\mathcal{K}$ holds. ∎
It was Schalk who proved that for a locally compact sober space $X$, the upper Vietoris topology and the Scott topology on $Q(X)$ coincide (see [20, Lemma 7.26]), meanwhile, [3, Proposition 5.4] reveals that each locally compact space is consonant. Therefore, combining Lemma 4.5 and Theorem 4.8, we further draw the conclusion in the following.
Corollary 4.9.
If $L$ is a locally compact sober dcpo, then $\mathcal{QH}(L)$ is isomorphic to $\mathcal{HQ}(L)$ under the function $\phi$ (or $\psi$).
5 The double power well-filtered dcpo
During the study of monads, there has always been a problem worth thinking about: Given two monads with underlying functors $S$ and $T$ on a category $\mathbf{C}$, does $T\circ S$ (or $S\circ T$) still carry the structure of a monad? Until 1969, Beck introduced the concepts of distributive laws and showed that the existence of distributive laws provides a sufficient condition for such a composition to form a monad [2].
Definition 5.1.
([2]) Let $\mathcal{S}=(\mathcal{S},\eta^{\mathcal{S}},\mu^{\mathcal{S}})$ and $\mathcal{T}=(\mathcal{T},\eta^{\mathcal{T}},\mu^{\mathcal{T}})$ be two monads on the category $\mathbf{C}$. A distributive law of $\mathcal{S}$ over $\mathcal{T}$ is a natural transformation $l:\mathcal{TS}\rightarrow\mathcal{ST}$ such that the following diagrams commute (Figure 1).
$\mathcal{T}$$\mathcal{TS}$$\mathcal{T}\eta^{\mathcal{S}}$$\mathcal{ST}$$\eta^{\mathcal{S}}_{\mathcal{T}}$$l$$\mathcal{S}$$\mathcal{TS}$$\mathcal{ST}$$\mathcal{S}\eta^{\mathcal{T}}$$\eta^{\mathcal{T}}_{\mathcal{S}}$$l$$\mathcal{TSS}$$\mathcal{STS}$$\mathcal{SST}$$l_{\mathcal{S}}$$\mathcal{S}l$$\mathcal{TS}$$\mathcal{ST}$$T\mu^{\mathcal{S}}$$\mu^{\mathcal{S}}_{\mathcal{T}}$$l$$\mathcal{TTS}$$\mathcal{TST}$$\mathcal{STT}$$\mathcal{T}l$$l_{\mathcal{T}}$$\mathcal{TS}$$\mathcal{ST}$$\mu^{\mathcal{T}}_{\mathcal{S}}$$\mathcal{S}\mu^{\mathcal{T}}$$l$
Figure 1
Definition 5.2.
([1]) Let $\mathcal{T}=(\mathcal{T},\eta,\mu)$ and $\mathcal{T}^{\prime}=(\mathcal{T}^{\prime},\eta^{\prime},\mu^{\prime})$ be two monads. A morphism of monads is a natural transformation $\lambda:\mathcal{T}\rightarrow\mathcal{T}^{\prime}$ such that the following two diagrams commute (Figure 2).
$Id$$\mathcal{T}$$\mathcal{T}^{\prime}$$\eta$$\eta^{\prime}$$\lambda$$\mathcal{TT}$$\mathcal{T^{\prime}T}$$\mathcal{T^{\prime}T^{\prime}}$$\mathcal{T}$$\mathcal{T}^{\prime}$$\lambda_{\mathcal{T}}$$\mathcal{T}^{\prime}\lambda$$\mu$$\mu^{\prime}$$\lambda$
Figure 2
Particularly, the morphisms of monads are called triple maps by Beck in [2]. Meanwhile, Beck also proved the following theorem.
Theorem 5.3.
([2]) Let $\mathcal{S}=(\mathcal{S},\eta^{\mathcal{S}},\mu^{\mathcal{S}})$, $\mathcal{T}=(\mathcal{T},\eta^{\mathcal{T}},\mu^{\mathcal{T}})$ be two monads on the category $\mathbf{C}$. Then the following statements are equivalent:
  1. 1.
    There are distributive laws $d:\mathcal{ST}\rightarrow\mathcal{TS};$
  2. 2.
    There are multiplications $\mu:\mathcal{TSTS}\rightarrow\mathcal{TS}$ such that
    • (a)
      $(\mathcal{TS},\eta^{\mathcal{T}}\eta^{\mathcal{S}},\mu)$ is a monad;
    • (b)
      The natural transformations $\eta^{\mathcal{T}}_{\mathcal{S}}:\mathcal{S}\rightarrow\mathcal{TS}$ and $\mathcal{T}\eta^{\mathcal{S}}:\mathcal{T}\rightarrow\mathcal{TS}$ are morphisms of monads.
Proposition 5.4.
$\phi$ and $\psi$ given in Definition 4.1 are distributive laws of $\mathcal{Q}$ over $\mathcal{H}$ and of $\mathcal{H}$ over $\mathcal{Q}$, respectively.
Proof.
It is a routine to check by their definitions. ∎
Based on the above results, we conclude that both $\mathcal{H}\mathcal{Q}$ and $\mathcal{Q}\mathcal{H}$ are monads on $\mathbf{WF}$. Naturally, the problem that needs further consideration is whether the $\mathcal{Q}\mathcal{H}$-algebras (or $\mathcal{H}\mathcal{Q}$-algebras) are related to the $\mathcal{Q}$-algebras and the $\mathcal{H}$-algebras. In the following, we will take the monad $\mathcal{Q}\mathcal{H}$ as the representative for an investigation.
We first give the unit $\gamma$ and the multiplication $\rho$ of $\mathcal{Q}\mathcal{H}$ with the following diagrams, which are closely related to that of $\mathcal{Q}$ and $\mathcal{H}$ according to Beck [2].
$\gamma:$$Id$$\mathcal{Q}$$\mathcal{H}$$\mathcal{Q}\mathcal{H}$$\theta$$\eta$$\mathcal{Q}\eta$$\theta_{\mathcal{H}}$$\rho:$$\mathcal{Q}\mathcal{H}\mathcal{Q}\mathcal{H}$$\mathcal{Q}\mathcal{Q}\mathcal{H}\mathcal{H}$$\mathcal{Q}\mathcal{H}\mathcal{H}$$\mathcal{Q}\mathcal{Q}\mathcal{H}$$\mathcal{Q}\mathcal{H}$$\mathcal{Q}\phi_{\mathcal{H}}$$\iota_{\mathcal{H}\mathcal{H}}$$\mathcal{Q}\mathcal{Q}\mu$$\mathcal{Q}\mu$$\iota_{\mathcal{H}}$
[…]
Definition 5.5.
([8]) Let $\mathbf{C}$ be a poset-enriched category. A monad $\mathcal{T}=(\mathcal{T},\eta,\mu)$ on $\mathbf{C}$ is called a right $\rm{KZ}$-monad if $\mathcal{T}$ is a poset-functor and subjects to $\eta_{TX}\leq T\eta_{X}$ for all $X\in\mathbf{C}$.
Left $\rm{KZ}$-monads are defined poset-dually, by reversing the inequality.
Remark 5.6.
[…]
  1. 1.
    For all $C\in\mathbf{C}$, an arrow $\alpha:\mathcal{T}C\rightarrow C$ is a structure map iff $\eta_{C}\dashv\alpha$ and $\alpha\circ\eta_{C}=id_{C}$.
  2. 2.
    $\mu_{C}\dashv\mathcal{T}\eta_{C}$ for all $C\in\mathbf{C}$.
Theorem 5.7.
([8]) Let $\mathcal{T}=(\mathcal{T},\eta,\mu)$ be a right $\rm{KZ}$-monad in a poset-enriched category $\mathbf{C}$. Then the following statements are equivalent for any object $C\in\mathbf{C}$.
  1. 1.
    $C$ is a $T$-algebra.
  2. 2.
    $C$ is injective over the right $T$-embeddings.
Note that a right $T$-embedding mentioned in the above theorem is an arrow in $\mathbf{C}$ satisfying that $Tf$ has a left adjoint $h$ (i.e., $h\circ Tf\leq id$ and $Tf\circ h\geq id$), and $h\circ Tf=id$.
Lemma 5.8.
The category $\mathbf{WF}$ is a poset-enriched category; $\mathcal{H}$ and $\mathcal{Q}$ are both poset-functors.
Proposition 5.9.
$\mathcal{H}$ is a left $\rm{KZ}$-monad and $\mathcal{Q}$ is a right $\rm{KZ}$-monad.
Proof.
The definitions of the two $\rm{KZ}$-monads make the verification trivial. ∎
Lemma 5.10.
([7]) Let $\mathcal{S}=(\mathcal{S},\eta_{\mathcal{S}},\mu_{\mathcal{S}})$ and $\mathcal{T}=(\mathcal{T},\eta_{\mathcal{T}},\mu_{\mathcal{T}})$ be two monads on a category $\mathbf{C}$ and $\lambda:\mathcal{S}\rightarrow\mathcal{T}$ a morphism of monads. Then every $\mathcal{T}$-algebra $(A,\alpha)$ can be equipped with an $\mathcal{S}$-algebra via $(A,\alpha)\mapsto(A,\alpha\circ\lambda)$. Moreover, a $\mathcal{T}$-algebra homomorphism $f:(A,\alpha)\rightarrow(B,\beta)$ induces an $\mathcal{S}$-algebra homomorphism $f:(A,\alpha\circ\lambda)\rightarrow(B,\beta\circ\lambda)$ in a functorial way.
Proposition 5.11.
Let $(L,\alpha)$ be a $\mathcal{Q}\mathcal{H}$-algebra in $\mathbf{WF}$. Then
  1. 1.
    $L$ is a $\mathcal{Q}$-algebra and an $\mathcal{H}$-algebra.
  2. 2.
    $\alpha$ is a $\mathcal{Q}$-algebra homomorphism and $\alpha\circ\phi$ is an $\mathcal{H}$-algebra homomorphism.
  3. 3.
    $\alpha$ is the unique $\mathcal{Q}\mathcal{H}$-structure map of $L$.
Proof.
(1): To prove that $L$ is a $\mathcal{Q}$-algebra, let $\alpha^{\mathcal{Q}}=\alpha\circ\mathcal{Q}\eta_{L}:\mathcal{Q}(L)\rightarrow L$. By Lemma 5.10, we know that $\alpha^{\mathcal{Q}}$ will be the corresponding structure map as long as $\mathcal{Q}\eta$ is a morphism of monads from $\mathcal{Q}$ to $\mathcal{QH}$. As it happens, it is Theorem 5.3 together with Proposition 5.4 that guarantees this fact. So $(L,\alpha^{\mathcal{Q}})$ is a $\mathcal{Q}$-algebra.
We proceed by defining $\alpha^{\mathcal{H}}=\alpha\circ\theta_{\mathcal{H}(L)}:\mathcal{H}(L)\rightarrow L$. Utilizing Theorem 5.3 and Proposition 5.4 again, we can get that $\theta_{\mathcal{H}}$ is a morphism of monads from $\mathcal{H}$ to $\mathcal{QH}$. So we conclude that $(L,\alpha^{\mathcal{H}})$ is an $\mathcal{H}$-algebra.
(2): By the fact that for each object $P$ in $\mathbf{WF}$, $(\mathcal{Q}(P),\iota_{P})$ is a $\mathcal{Q}$-algebra, we know $(\mathcal{QH}(L),\iota_{\mathcal{H}(L)})$ is also a $\mathcal{Q}$-algebra. Then according to the definition of $\mathcal{Q}$-algebra homomorphism, we need to show the diagram
$\mathcal{Q}^{2}\mathcal{H}(L)$$\mathcal{QH}(L)$$\mathcal{Q}(L)$$L$$\iota_{\mathcal{H}(L)}$$\alpha^{\mathcal{Q}}$$\mathcal{Q}\alpha$$\alpha$
commutes, that is, $\alpha^{\mathcal{Q}}\circ\mathcal{Q}\alpha=\alpha\circ\iota_{\mathcal{H}(L)}$. Note that
$\alpha^{\mathcal{Q}}\circ\mathcal{Q}\alpha$ $=\alpha\circ\mathcal{Q}\eta_{L}\circ\mathcal{Q}\alpha$
$=\alpha\circ\mathcal{Q}(\eta_{L}\circ\alpha)$
$=\alpha\circ\mathcal{Q}(\mathcal{H}\alpha\circ\eta_{\mathcal{QH}(L)})\quad\mathrm{by\ naturality\ of\ \eta}$
$=\alpha\circ\mathcal{QH}\alpha\circ\mathcal{Q}\eta_{\mathcal{QH}(L)}$
$=\alpha\circ\rho_{L}\circ\mathcal{Q}\eta_{\mathcal{QH}(L)}\qquad\mathrm{since\ (L,\alpha)\ is\ a\ \mathcal{QH}{-}algebra},$
which means that we just need to show
$\rho_{L}\circ\mathcal{Q}\eta_{\mathcal{QH}}=\iota_{\mathcal{H}(L)}:\mathcal{Q}^{2}\mathcal{H}(L)\rightarrow\mathcal{QH}(L)$.
Assume $\mathbf{K}\in\mathcal{Q}^{2}\mathcal{H}(L)$. One can calculate that $\iota_{\mathcal{H}(L)}(\mathbf{K})=\bigcup\mathbf{K}$ and
$\rho_{L}\circ\mathcal{Q}\eta_{\mathcal{QH}}(\mathbf{K})={\uparrow}\{\bigcup\mathcal{C}:\mathcal{C}\in\bigcup{\uparrow}\{\phi_{\mathcal{H}}(\mathbf{A}):\mathbf{A}\in{\uparrow}\{{\downarrow}\{\mathcal{K}\}:\mathcal{K}\in\mathbf{K}\}\}\}$,
in which we take $\rho=\mathcal{Q}\mu\circ\rho_{\mathcal{HH}}\circ\mathcal{Q}\phi_{\mathcal{H}}$. On the one hand, for any $C\in\bigcup\mathbf{K}$, $C\in\mathcal{K}_{0}$ for some $\mathcal{K}_{0}\in\mathbf{K}$, which means ${\downarrow}\{C\}\bigcap\mathcal{K}_{0}\neq\emptyset$, thus for each $\mathcal{M}\supseteq\mathcal{K}_{0}$, i.e., $\mathcal{M}\in{\downarrow}\{\mathcal{K}_{0}\}$, we have ${\downarrow}\{C\}\bigcap\mathcal{M}\neq\emptyset$. It follows that ${\downarrow}\{C\}\in\phi_{\mathcal{H}}({\downarrow}\{\mathcal{K}_{0}\})$, and further, ${\downarrow}\{C\}\in\bigcup{\uparrow}\{\phi_{\mathcal{H}}(\mathbf{A}):\mathbf{A}\in{\uparrow}\{{\downarrow}\{\mathcal{K}\}:\mathcal{K}\in\mathbf{K}\}\}$. Then the fact that $C=\bigcup{\downarrow}\{C\}$ implies $C\in\rho_{L}\circ\mathcal{Q}\eta_{\mathcal{QH}}(\mathbf{K})$. So $\iota_{\mathcal{H}(L)}(\mathbf{K})\subseteq\rho_{L}\circ\mathcal{Q}\eta_{\mathcal{QH}}(\mathbf{K})$ holds. On the other hand, for each $\mathcal{C}\in\bigcup{\uparrow}\{\phi_{\mathcal{H}}(\mathbf{A}):\mathbf{A}\in{\uparrow}\{{\downarrow}\{\mathcal{K}\}:\mathcal{K}\in\mathbf{K}\}\}$, there exist an $\mathbf{M}\in\mathcal{QHH}(L)$ and an $\mathbf{A}_{0}\in{\uparrow}\{{\downarrow}\{\mathcal{K}\}:\mathcal{K}\in\mathbf{K}\}$ such that $\mathcal{C}\in\mathbf{M}\subseteq\phi_{\mathcal{H}}(\mathbf{A}_{0})$. Note that $\mathbf{A}_{0}\in{\uparrow}\{{\downarrow}\{\mathcal{K}\}:\mathcal{K}\in\mathbf{K}\}$ means that $\mathcal{K}_{0}\in\mathbf{A}_{0}$ for some $\mathcal{K}_{0}\in\mathbf{K}$, thus $\mathcal{C}\bigcap\mathcal{K}_{0}\neq\emptyset$. Suppose that $C_{0}$ is in this intersection. Then $C_{0}\subseteq\bigcup\mathcal{C}$; moreover, by the saturation of $\mathcal{K}_{0}$, $\bigcup\mathcal{C}\in\mathcal{K}_{0}$, which implicates $\bigcup\mathcal{C}\in\bigcup\mathbf{K}$. Hence, $\rho_{L}\circ\mathcal{Q}\eta_{\mathcal{QH}}(\mathbf{K})\subseteq\iota_{\mathcal{H}(L)}(\mathbf{K})$.
As argued above, we have $\rho_{L}\circ\mathcal{Q}\eta_{\mathcal{QH}}(\mathbf{K})=\iota_{\mathcal{H}(L)}(\mathbf{K})$ and conclude that $\alpha$ is a $\mathcal{Q}$-algebra homomorphism.
As for $\alpha\circ\phi$, since $(L,\alpha^{\mathcal{H}})$ and $(\mathcal{HQ}(L),\mu_{\mathcal{Q}(L)})$ are both $\mathcal{H}$-algebras and by the definition of $\mathcal{H}$-algebra homomorphism, we need to verify that the diagram below commutes,
$\mathcal{H}^{2}\mathcal{Q}(L)$$\mathcal{HQ}(L)$$\mathcal{H}(L)$$L$$\mu_{\mathcal{Q}(L)}$$\alpha^{\mathcal{H}}$$\mathcal{H}(\alpha\circ\phi)$$\alpha\circ\phi$
more precisely, $\alpha^{\mathcal{H}}\circ\mathcal{H}(\alpha\circ\phi)=\alpha\circ\phi\circ\mu_{\mathcal{Q}(L)}$ holds. One can see that
$\alpha^{\mathcal{H}}\circ\mathcal{H}(\alpha\circ\phi)$ $=\alpha\circ\theta_{\mathcal{H}(L)}\circ\mathcal{H}\alpha\circ\mathcal{H}\phi$
$=\alpha\circ\mathcal{QH}\alpha\circ\theta_{\mathcal{HQH(L)}}\circ\mathcal{H}\phi\quad\mathrm{by\ naturality\ of\ \theta}$
$=\alpha\circ\rho_{L}\circ\theta_{\mathcal{HQH(L)}}\circ\mathcal{H}\phi.$
Thus it is enough to show $\rho_{L}\circ\theta_{\mathcal{HQH(L)}}\circ\mathcal{H}\phi=\phi\circ\mu_{\mathcal{Q}(L)}$. The verification process is a routine similar to that of $\rho_{L}\circ\mathcal{Q}\eta_{\mathcal{QH}}=\iota_{\mathcal{H}(L)}$. Hence, we could draw the conclusion that $\alpha\circ\phi$ is an $\mathcal{H}$-algebra homomorphism.
[…]
$k({\uparrow}\{A\})=k\circ\theta_{\mathcal{H}(L)}(A)=m\circ\theta_{\mathcal{H}(L)}(A)=m({\uparrow}\{A\})$,
which induces
$k(\mathcal{K})=k(\bigcup_{A\in\mathcal{K}}{\uparrow}\{A\})=\bigcup_{A\in\mathcal{K}}k({\uparrow}\{A\})=\bigcup_{A\in\mathcal{K}}m({\uparrow}\{A\})=m(\bigcup_{A\in\mathcal{K}}{\uparrow}\{A\})=m(\mathcal{K})$
for each $\mathcal{K}\in\mathcal{QH}(L)$, that is to say, $k=m$. Note that $\alpha\circ\theta_{\mathcal{H}(L)}=\alpha^{\mathcal{H}}=\beta^{\mathcal{H}}=\beta\circ\theta_{\mathcal{H}(L)}$. Then the uniqueness of $m$ enables us to conclude that $\alpha=\beta$.
Proposition 5.12.
If the distributive law $\phi$ is an isomorphism and $L$ is a $\mathcal{QH}$-algebra, then $L$ is a frame.
Proof.
The fact given in Proposition 5.11 that $L$ is a $\mathcal{Q}$-algebra and an $\mathcal{H}$-algebra yields that $L$ is indeed a meet continuous complete lattice by Corollary 3.8 and Proposition 3.16. Thus the proof will be completed if we show that $L$ is distributive.
$\mathbf{Claim~{}1}$: $\mathcal{QH}(L)$ is a distributive lattice.
Obviously, the union of arbitrary finite elements in $\mathcal{QH}(L)$ is still compact saturated. By Lemma 4.2, we know the complete lattice $\mathcal{H}(L)$ endowed with the Scott topology is coherent, that is, the arbitrary finite intersection of compact saturated subsets of $\mathcal{H}(L)$ is still in $\mathcal{QH}(L)$. Thus $\mathcal{QH}(L)$ is a lattice, in which the supremum and infimum of each finite subset are their union and intersection, respectively. This contributes to the distributivity of $\mathcal{QH}(L)$.
$\mathbf{Claim~{}2}$: The structure map $\alpha:\mathcal{QH}(L)\rightarrow L$ is a lattice homomorphism.
It follows from Proposition 5.11 that $\alpha$ is a $\mathcal{Q}$-homomorphism. Thus $\alpha$ preserves finite infs by Lemma 3.15. For any $\mathcal{K}_{1},\mathcal{K}_{2}\in\mathcal{QH}(L)$, since $\phi$ is surjective, there exist $\mathcal{A}_{1},\mathcal{A}_{2}\in\mathcal{HQ}(L)$ such that $\phi(\mathcal{A}_{i})=\mathcal{K}_{i},i=1,2$. Thus
$\alpha(\mathcal{K}_{1}\vee\mathcal{K}_{2})$ $=\alpha(\mathcal{K}_{1}\cap\mathcal{K}_{2})=\alpha(\phi(\mathcal{A}_{1})\cap\phi(\mathcal{A}_{2}))=\alpha(\phi(\mathcal{A}_{1}\cup\mathcal{A}_{2}))$
$=\alpha\circ\phi(\mathcal{A}_{1}\vee\mathcal{A}_{2})=\alpha\circ\phi(\mathcal{A}_{1})\vee\alpha\circ\phi(\mathcal{A}_{2})$
$=\alpha(\mathcal{K}_{1})\vee\alpha(\mathcal{K}_{2}),$
where the penultimate equation holds thanks to the fact that $\alpha\circ\phi$ is an $\mathcal{H}$-homomorphism proved in Proposition 5.11.
Now we consider the functions $\alpha\circ\varphi_{i}\circ(\theta_{\mathcal{H}}\circ\eta_{L})^{3}:L\times L\times L\rightarrow L,\ i=1,2$, where $(\theta_{\mathcal{H}}\circ\eta_{L})^{3}:L\times L\times L\rightarrow\mathcal{QH}(L)\times\mathcal{QH}(L)\times\mathcal{QH}(L)$ is defined as
$(\theta_{\mathcal{H}}\circ\eta_{L})^{3}(x,y,z)=({\uparrow}\{{\downarrow}x\},{\uparrow}\{{\downarrow}y\},{\uparrow}\{{\downarrow}z\})$, for any $x,y,z\in L$,
and $\varphi_{i}:(\mathcal{QH}(L))^{3}\rightarrow\mathcal{QH}(L),\ i=1,2$ are defined below: for each triple $(\mathcal{K}_{1},\mathcal{K}_{2},\mathcal{K}_{3})\in(\mathcal{QH}(L))^{3}$,
$\varphi_{1}(\mathcal{K}_{1},\mathcal{K}_{2},\mathcal{K}_{3})=\mathcal{K}_{1}\wedge(\mathcal{K}_{2}\vee\mathcal{K}_{3})$,
$\varphi_{2}(\mathcal{K}_{1},\mathcal{K}_{2},\mathcal{K}_{3})=(\mathcal{K}_{1}\wedge\mathcal{K}_{2})\vee(\mathcal{K}_{1}\wedge\mathcal{K}_{3})$.
One can calculate that $\alpha\circ\varphi_{1}\circ(\theta_{\mathcal{H}}\circ\eta_{L})^{3}(x,y,z)=x\wedge(y\vee z)$ and $\alpha\circ\varphi_{2}\circ(\theta_{\mathcal{H}}\circ\eta_{L})^{3}(x,y,z)=(x\wedge y)\vee(x\wedge z)$. It is obvious that $\varphi_{1}=\varphi_{2}$ by Claim 1. Thus $x\wedge(y\vee z)=(x\wedge y)\vee(x\wedge z)$, that is, $L$ satisfies the distributive law. In conclusion, $L$ is a frame. ∎
Proposition 5.13.
Every $\mathcal{QH}$-algebra homomorphism is a frame homomorphism.
Proof.
Let $f:(L,\alpha)\rightarrow(M,\beta)$ be a $\mathcal{QH}$-algebra homomorphism. It was shown in Theorem 5.3 that $\mathcal{Q}\eta:\mathcal{Q}\rightarrow\mathcal{QH}$ and $\theta_{\mathcal{H}}:\mathcal{H}\rightarrow\mathcal{QH}$ are morphisms of monads; hence, by Lemma 5.10, we know $f:(L,\alpha\circ\mathcal{Q}\eta_{L})\rightarrow(M,\beta\circ\mathcal{Q}\eta_{M})$ is a $\mathcal{Q}$-algebra homomorphism and $f:(L,\alpha\circ\theta_{\mathcal{H}(L)})\rightarrow(M,\beta\circ\theta_{\mathcal{H}(L)})$ is an $\mathcal{H}$-algebra homomorphism. It follows that $f$ preserves finite infs and arbitrary sups, that is to say, $f$ is a frame homomorphism.
Let $\mathbf{Frm}$ be the category of all frames and all frame homomorphisms.
Theorem 5.14.
If the distributive law $\phi$ is an isomorphism, then the Eilenberg-Moore category of $\mathcal{QH}$ over $\mathbf{WF}$ is a subcategory of the category $\mathbf{Frm}$.
Remark 5.15.
See the deflationary semilattice in $\mathbf{WF}$ quoted in Example 3.17 $\mathcal{L}$ again, it has been proven not to be a $\mathcal{Q}$-algebra. Thus evidenced by Proposition 5.11, $\mathcal{L}$ is not a $\mathcal{QH}$-algebra, which indicates that $\mathcal{QH}$ can not give the free deflationary semilattices for $\mathbf{WF}$.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.