### v3

# Commutation of Smyth and Hoare Power Constructions in Well-filtered Dcpos

Abstract
Prior work [11] established a commutativity result for the Hoare power construction and a modified version of the Smyth power construction consisting of strongly compact sets, which is defined for $\mathcal{U}_{S}$-admitting dcpos, where $\mathcal{U}_{S}$-admissability is well-filteredness with compact sets replaced by strongly compact sets. In this paper, we consider the Hoare power construction $\mathcal{H}$ and the Smyth power construction $\mathcal{Q}$ on the category $\mathbf{WF}$ of well-filtered dcpos with Scott-continuous maps. Actually, the functors $\mathcal{H}$ and $\mathcal{Q}$ can be extended to monads. We prove that $\mathcal{H}$ and $\mathcal{Q}$ commute, that is, $\mathcal{HQ}(L)$ is isomorphic to $\mathcal{QH}(L)$ for a well-filtered dcpo $L$, if and only if $L$ satisfies a property similar to consonance that we call $(\mathrm{KC})$ and the Scott topology coincides with the upper Vietoris topology on $\mathcal{Q}(L)$. We also investigate the Eilenberg-Moore category of the monad composed by $\mathcal{H}$ and $\mathcal{Q}$ under a distributive law $\phi$ on $\mathbf{WF}$ and characterize it to be a subcategory of the category $\mathbf{Frm}$, which is composed of all frames and all frame homomorphisms.
keywords:
Well-filtered dcpos , Power constructions , Monad , $\mathrm{Eilenberg}$-$\mathrm{Moore}$ category , Commutativity
MSC:
18C15, 18C20
t1t1footnotetext: This work is supported by the National Natural Science Foundation of China (No.12231007)
1 Introduction
Non-deterministic choices are important semantic concepts that offer new insights in designing more powerful programming languages. They were originally modeled by the power constructions on the category of domains. The most common of these are the Hoare and the Smyth powerdomains [22], which can respectively serve as the denotational semantics for angelic and demonic non-deterministic choices. Notably, the Hoare and the Smyth powerdomains also give rise to monads. In 1991, Eugenio Moggi showed that, within the category of sets, the collection of all finite sets could serve as denotations for nondeterministic programming languages [19]. He further proved that this structure forms a monad on the category of sets, thereby linking power constructions with monads. Over time, monads have evolved into a fundamental tool for constructing denotational semantic models in functional programming languages.
In 1969, Jon Beck proposed the notion of the distributive law between two monads defined on the same category [2]. Such a distributive law ensures that the composition of these two monads remains a monad, providing a new approach to establishing denotational semantic models in functional programming languages. Moreover, the two composite monads obtained under different composition orders may differ; if they coincide, i.e., the monad functors induce isomorphic mappings on objects, then the monads are said to obey a commutative law. Such commutativity ensures that the order of program execution becomes immaterial in computation. This practical significance raises two fundamental problems: (1) Across which categories does a distributive law exist between the two denotational semantic models for nondeterministic programming languages the Hoare power monad and the Smyth power monad? (2) Within a specified category, under what conditions do these two monads commute?
In domain theory, numerous scholars have conducted in-depth investigations into the above two problems. In 1990, Kevin Flannery and Johannes Martin employed information systems to demonstrate the commutativity between the Hoare power monad and the Smyth power monad in the category of bounded complete algebraic domains [6]. Subsequently, Reinhold Heckmann used an entirely distinct method to prove that for any full subcategory of the category of dcpos, the commutativity of Hoare and Smyth power monads is guaranteed for variants of the two monads that are defined as the free inflationary semilattice and the free deflationary semilattice, respectively [12]. Within the category of domains, the Hoare powerdomain induces the free upper semilattices, while the Smyth powerdomain yields the free meet-continuous semilattices [9]. Here, the upper semilattices are actually inflationary semilattices, and the meet-continuous semilattices are deflationary semilattices. Thus the Hoare power monad and the Smyth power monad commute in the category of domains.
Note that many categories have the category of domains as a full subcategory. The versions of Hoare and Smyth power monads considered in [12], i.e. defined as free constructions, are well-defined for all dcpos and do commute on the category of all dcpos. However, the specific free constructions here are not clear. So the versions of the power monads defined via closed and saturated compact subsets are different from the free constructions in general. It implies that the restriction to well-filtered dcpos (called $\mathcal{U}_{K}$-admitting in [11]) is necessary to obtain a well-behaved Smyth contruction. This issue has been evolving continuously. Concretely, Heckmann in [11] investigated the commutativity between the Hoare power construction $\mathcal{L}$ and a modified Smyth power construction $\mathcal{U}_{S}$ on the category of $\mathcal{U}_{S}$-admitting dcpos [11]. In this work, the extended Smyth power construction is determined by all strongly compact subsets, not the traditional Smyth power construction based on compact subsets. Thus, the $\mathcal{U}_{S}$-admissibility is strictly weaker than well-filteredness. Heckmann introduced the property of being $\mathcal{U}_{S}$-conformal and showed that for a dcpo $L$, $L$ is $\mathcal{U}_{S}$-conformal if and only if $\mathcal{L}(\mathcal{U}_{S}(L))$ is isomorphic to $\mathcal{U}_{S}(\mathcal{L}(L))$, i.e., established the sufficient and necessary conditions for the commutativity. Meanwhile, they also pointed out that the corresponding commutativity could not be characterized in the category of well-filtered dcpos.
In this paper, we focus on the category of well-filtered dcpos and investigate the commutativity between the associated Hoare and Smyth power monads. In Section 3, we define the Hoare and Smyth constructions for a well-filtered dcpo $L$ as follows: The Hoare construction $\mathcal{H}$ is the set of all Scott closed subsets of $L$, ordered by inclusion; the Smyth construction $\mathcal{Q}$ is the set of all Scott compact saturated subsets of $L$, ordered by reverse inclusion order. We explicitly show that $\mathcal{Q}$ does not yield a free construction for the category $\mathbf{WF}$ of well-filtered dcpos with Scott-continuous maps, which confirms the significance and necessity of our research. In Section 4, we introduce a property called $(\mathrm{KC})$, which is weaker than the consonance in terms of posets. Based on this we proved that $\mathcal{H}$ and $\mathcal{Q}$ commute in the sense of that $\mathcal{HQ}(L)$ is isomorphic to $\mathcal{QH}(L)$ under a certain function, if and only if $L$ satisfies the property $(\mathrm{KC})$ and the Scott topology coincides with the upper Vietoris topology on $\mathcal{Q}(L)$. It is necessary to point out that in [4], de Brecht and Kawai also provided a Smyth construction via compact saturated subsets and investigated the commutativity between the Smyth construction and the Hoare construction. However, our work differs from theirs, since they consider the category of quasi-polish topological spaces while we explore that of well-filtered dcpos.
Notably, unlike Heckmann’s modified Smyth power construction $\mathcal{U}_{S}$, $\mathcal{Q}$ in this paper induces the Smyth power monad on $\mathbf{WF}$. This relies crucially on Xu’s result [26] that the Smyth power space endowed with the Scott topology of any well-filtered space remains well-filtered. Furthermore, since every complete lattice with the Scott topology is well-filtered [25, Corollary 3.2], we conclude that $\mathcal{H}$ also forms a monad on $\mathbf{WF}$. Thus in Section 5, we analyze the properties of the composite monad. The morphism $\phi$ defined in Definition 4.1 is shown to be a distributive law, which ensures that the composition $\mathcal{QH}$ remains a monad on $\mathbf{WF}$. Finally, motivated by the growing interest in Eilenberg-Moore algebras, we characterize the Eilenberg-Moore category of $\mathcal{QH}$ as a subcategory of Frm the category of frames and frame homomorphisms.
2 Preliminaries
In this section, we will give a brief introduction to some basic concepts and notions to be used in this paper. One can refer to [1, 9, 10] for more details.
Given a poset $P$, $A\subseteq P$ is an upper set if $A={\uparrow}A$, where ${\uparrow}A=\{x\in P:x\geq a\ \mathrm{for\ some}\ a\in A\}$. The lower set is defined dually. The subset $D$ of $P$ is directed if every finite subset of $D$ has an upper bound in $D$. If each directed subset in a poset $P$ has a supremum, we call $P$ is a directed complete partially ordered set (dcpo). A subset $U$ of $P$ is called Scott open iff it is an upper set and for each directed subset $D$ whose supremum exists, written as $\sup D$, $\sup D\in U$ implies $D\cap U\neq\emptyset$. The complement of a Scott open set is called Scott closed. The collection of all Scott open subsets, denoted by $\sigma(P)$, forms a topology called Scott topology and in general, we write $\Sigma P=(P,\sigma(P))$. Besides, we use $\Gamma(P)$ to denote the set of all Scott closed subsets of $P$. A continuous map between two posets equipped with Scott topology is called Scott-continuous. Generally, $\mathbf{DCPO}$ is used to denote the category of all dcpos with Scott-continuous mappings.
Let $X$ be a topological space. We use $\mathcal{O}(X)$ denote the collection consisting of all open subsets of $X$. When $X$ is a $T_{0}$ space, the specialization preorder $\leq$ defined by $x\leq y$ iff $x\in cl(\{y\})$ is a partial order and is called specialization order. A subset of $X$ is saturated if it is the intersection of open sets, or equivalently, it is an upper set in the order of specialization. $X$ is called well-filtered if for each filter basis $\mathcal{C}$ of compact saturated subsets and any open subset $U\in\mathcal{O}(X)$ with $\bigcap\mathcal{C}\subseteq U$, $K\subseteq U$ for some $K\in\mathcal{C}$. We usually say a dcpo $P$ is well-filtered if it equipped with the Scott topology is a well-filtered space. Furthermore, a space $X$ is called coherent if the intersection of any two compact saturated sets of $X$ is again compact.
In this paper, we denote the set of all compact saturated subsets (including the empty set) and that of all closed subsets of a space $X$ by $\mathcal{Q}(X)$ and $\Gamma(X)$, respectively. The topology on $\mathcal{Q}(X)$ generated by $\{\square U:U\in\mathcal{O}(X)\}$ as a base, where $\Box U=\{K\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{Q}(X)}:K\subseteq U\}$, is called the upper Vietoris topology and denoted by $\upsilon({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{Q}(X)})$. The upper topology on $\Gamma(X)$ coincides with the lower Vietoris topology, which is written as $\nu(\Gamma(X))$ and generated by $\{\lozenge U:U\in\mathcal{O}(X)\}$ as a subbase, where $\lozenge U=\{A\in\Gamma(X):A\cap U\neq\emptyset\}$.
Definition 2.1.
[…]
Definition 2.2.
([1]) A monad $\mathcal{T}$ in a category $\mathbf{C}$ is a triple consisting of an endofunctor $\mathcal{T}:\mathbf{C}\rightarrow\mathbf{C}$ and two natural transformations: the unit $\eta:Id_{\mathbf{C}}\rightarrow\mathcal{T}$ and the multiplication $\mu:\mathcal{T}^{2}\rightarrow\mathcal{T}$ such that the following diagrams commute.
$\mathcal{T}$$\mathcal{T}^{2}$$\mathcal{T}$$\mathcal{T}$$\eta_{\mathcal{T}}$$\mathcal{T}\eta$$\mu$$\mathcal{T}^{3}$$\mathcal{T}^{2}$$\mathcal{T}^{2}$$\mathcal{T}$$\mathcal{T}\mu$$\mu\mathcal{T}$$\mu$$\mu$
Definition 2.3.
([1]) Let $\mathcal{T}=(\mathcal{T},\eta,\mu)$ be a monad on the category $\mathbf{C}$.
  1. 1.
    A pair $(C,\alpha)$ is a $\mathcal{T}$-algebra (or Eilenberg-Moore algebra of $\mathcal{T}$) if $\alpha:\mathcal{T}C\rightarrow C$ is a morphism in $\mathbf{C}$ that satisfies:
    ($\mathbf{Associativity}$) $\alpha\circ\mu_{C}=\alpha\circ\mathcal{T}\alpha$,  ($\mathbf{Unit\;law}$) $\alpha\circ\eta_{C}=id_{C}$.
    In this case, $\alpha$ is called a structure map.
  2. 2.
    Let $(A,\alpha_{A})$ and $(B,\alpha_{B})$ be $\mathcal{T}$-algebras. A $\mathcal{T}$-algebra homomorphism from $(A,\alpha_{A})$ to $(B,\alpha_{B})$ is an arrow $h:A\rightarrow B$ satisfying:
    $(\mathbf{Homomorphism\;law})\ h\circ\alpha_{A}=\alpha_{B}\circ\mathcal{T}h$.
The category composed of $\mathcal{T}$-algebras and their homomorphisms is called Eilenberg-Moore category of $\mathcal{T}$ over $\mathbf{C}$.
Proposition 2.4.
Let $\mathcal{T}=(\mathcal{T},\eta,\mu)$ be a monad on the category $\mathbf{C}$. Then the following properties hold:
  1. 1.
    For every object $C$ in $\mathbf{C}$, $(\mathcal{T}C,\mu_{C})$ is a $\mathcal{T}$-algebra.
  2. 2.
    If $(C,\alpha)$ is a $\mathcal{T}$-algebra, then $\alpha:(\mathcal{T}C,\mu_{C})\rightarrow(C,\alpha)$ is a $\mathcal{T}$-homomorphism.
  3. 3.
    If $h_{1},h_{2}:(\mathcal{T}C,\mu_{C})\rightarrow(C^{\prime},\alpha)$ are $T$-homomorphisms with $h_{1}\circ\eta_{C}=h_{2}\circ\eta_{C}$, then $h_{1}=h_{2}$.
Proof.
(1) and (2) are clear from the definitions, more precisely, one can refer to [18]. Part (3) holds because $h_{i}=h_{i}\circ\mu_{C}\circ\mathcal{T}\eta_{C}=\alpha\circ\mathcal{T}h_{i}\circ\mathcal{T}\eta_{C}=\alpha\circ\mathcal{T}(h_{i}\circ\eta_{C})$.
3 The Hoare and Smyth power constructions
3.1 The Hoare power construction
Definition 3.1.
The Hoare power construction $\mathcal{H}(L)$ for a well-filtered dcpo $L$ is the set of all Scott closed subsets of $L$ ordered by inclusion.
We now focus on constructing mutually adjoint functors, through which the composite functor induces the monad $\mathcal{H}$ on $\mathbf{WF}$. Clearly, for each well-filtered dcpo $L$, its Hoare power construction is a complete lattice. Let $\mathbf{CL}$ denote the category of complete lattices with morphisms preserving arbitrary suprema.
Consider the morphisms in $\mathbf{WF}$. For any Scott-continuous mapping $f:L\rightarrow M$ between well-filtered dcpos $L$ and $M$, we define $\mathcal{H}(f):\mathcal{H}(L)\rightarrow\mathcal{H}(M)$ as
$\mathcal{H}(f)(A)=cl(f(A))$, for any $A\in\mathcal{H}(L)$.
For any collection $\{A_{i}:i\in I\}\subseteq\mathcal{H}(L)$, since $\mathcal{H}(f)$ clearly is order preserving, $\sup_{i\in I}\mathcal{H}(f)(A_{i})\subseteq\mathcal{H}(f)(\sup_{i\in I}A_{i})$ holds. Meanwhile,
$\mathcal{H}(f)(\sup_{i\in I}A_{i})=\mathcal{H}(f)(cl(\bigcup_{i\in I}A_{i}))\subseteq cl(f(\bigcup_{i\in I}A_{i}))=cl(\bigcup_{i\in I}f(A_{i}))=\sup_{i\in I}\mathcal{H}(f)(A_{i})$,
where the second symbol $\subseteq$" is guaranteed by the continuity of $f$. Thus we conclude that $\mathcal{H}(f)$ preserves arbitrary suprema. Additionally, one can verify that $\mathcal{H}$ preserves the identities and the compositions of any two morphisms. Along with the above, we can obtain the following result.
Lemma 3.2.
$\mathcal{H}$ is a functor from $\mathbf{WF}$ to $\mathbf{CL}$.
It follows from [25, Corollary 3.2] that each complete lattice endowed with the Scott topology is well-filtered. So there is a forgetful functor from $\mathbf{CL}$ to $\mathbf{WF}$. Andrea Schalk has proven that the functor $P_{D}^{b}:\mathbf{DCPO}\rightarrow\mathbf{CL}$ is left adjoint to the forgetful functor $U:\mathbf{CL}\rightarrow\mathbf{DCPO}$ (see [21, Theorem 6.2]). Note that if the functor $P_{D}^{b}$ is restricted from the category $\mathbf{DCPO}$ to its full subcategory $\mathbf{WF}$, it is just the functor $\mathcal{H}$. This immediately yields the following result.
Proposition 3.3.
The functor $\mathcal{H}:\mathbf{WF}\rightarrow\mathbf{CL}$ is left adjoint to the forgetful functor $U:\mathbf{CL}\rightarrow\mathbf{WF}$.
Since each pair of adjunction determines a monad, [1, Proposition 4.2.1], we deduce that the triple $(U\circ\mathcal{H},\eta,U\varepsilon_{\mathcal{H}})$, where $\eta$ and $\varepsilon$ are the unit and counit respectively, turns into a monad on $\mathbf{WF}$. For ease of expressions, we will abbreviate $U\circ\mathcal{H}$ as $\mathcal{H}$. Meanwhile, when $U\varepsilon_{\mathcal{H}}$ works on a well-filtered $\mathrm{dcpo}$ $L$, one can calculate $U\varepsilon_{\mathcal{H}(L)}$ as the natural transformation from $\mathcal{H}(\mathcal{H}(L))$ to $\mathcal{H}(L)$ that maps $\mathcal{A}$ to $\sup_{\mathcal{H}(L)}\mathcal{A}$.
Lemma 3.4.
Let $L$ be a dcpo and $\mathcal{A}$ a Scott closed subset of $\mathcal{H}(L)$. Then $\bigcup\mathcal{A}$ is Scott closed.
Proof.
It is similar to that of [14, Proposition 2.2].
It follows from this lemma that $\sup_{\mathcal{H}(L)}\mathcal{A}=\bigcup\mathcal{A}$. Now we conclude that
Theorem 3.5.
The endofunctor $\mathcal{H}$ together with the unit $\eta:Id\rightarrow U\circ\mathcal{H}$ and the multiplication $\mu:\mathcal{H}^{2}\rightarrow\mathcal{H}$ forms a monad on $\mathbf{WF}$, where $\eta$ and $\mu$ are defined concretely for a well-filtered dcpo $L$ as follows:
$\eta_{L}(x)={\downarrow}x$, for any $x\in L$
and
$\mu_{L}(\mathcal{A})=\bigcup\mathcal{A}$, for any $\mathcal{A}\in\mathcal{H}(\mathcal{H}(L))$.
Combining the fact that $\mathbf{CL}$ is the category of unital inflationary semilattices for $\mathbf{WF}$ with that the Hoare power well-filtered dcpo has the universal property, we have that
Corollary 3.6.
$\mathcal{H}$ gives the free unital inflationary semilattice for $\mathbf{WF}$.
Lemma 3.7.
Let $L,M$ be well-filtered dcpos. Then
  1. 1.
    $(L,\alpha)$ is an $\mathcal{H}$-algebra if and only if $L$ is a complete lattice.
  2. 2.
    $f:(L,\alpha)\rightarrow(M,\beta)$ is an $\mathcal{H}$-algebra homomorphism if and only if $f$ is a complete lattice homomorphism.
Proof.
Based on their definitions, it is easy to verify them.
Corollary 3.8.
The Eilenberg-Moore category of $\mathcal{H}$ over $\mathbf{WF}$ is just $\mathbf{CL}$.
3.2 The Smyth power construction
Xu et al. proved that for a well-filtered space $X$, $\Sigma\mathcal{Q}(X)$ is still well-filtered [26, Theorem 6.5], which indicates that for each well-filtered dcpo $L$, $\mathcal{Q}(L)$ with the Scott topology is well-filtered, that is, $\mathcal{Q}(L)$ is an object in $\mathbf{WF}$. This fact makes us give a well-defined Smyth power construction.
Definition 3.9.
The Smyth power construction for a well-filtered dcpo is the set of all Scott compact saturated subsets ordered by reverse inclusion.
Recall that, inspired by the facts that for a sober space $X$, the set $\mathcal{Q}(X)$ ordered by reverse inclusion is isomorphic to the set of Scott open filters in $\mathcal{O}(X)$ ordered by inclusion, and for a dcpo $L$, the set of all compact saturated subsets $\mathcal{Q}(L)$ ordered by reverse inclusion is no longer a dcpo, Andrea Schalk focused on the set of all Scott open filters $OFilt(\sigma(L))$, proved that $OFilt(\sigma(L))$ ordered by inclusion remains a dcpo. Concerning this, she gave an endofunctor on the category $\mathbf{DCPO}$ and yields a monad (one can refer to [21, Lemma 7.5, Theorem 7.7]). Fortunately, for a well-filtered dcpo $L$, the set of all Scott compact saturated subsets ordered by reverse inclusion remains a well-filtered dcpo as shown in [26]. Next we will construct an endofunctor $\mathcal{Q}$ on $\mathbf{WF}$ and verify that it is also a monad.
Lemma 3.10.
[11, Theorem 3.9] Let $f:L\rightarrow M$ be a Scott-continuous mapping between well-filtered dcpos $L$ and $M$. Then $\mathcal{Q}(f):\mathcal{Q}(L)\rightarrow\mathcal{Q}(M)$ defined as
$\mathcal{Q}(f)(K)={\uparrow}f(K)$, for any $K\in\mathcal{Q}(L)$
is Scott-continuous.
One can easily verify that $\mathcal{Q}$ preserves the identities and the compositions of any two morphisms. So, $\mathcal{Q}$ is an endofunctor on $\mathbf{WF}$.
Lemma 3.11.
Let $L$ be a well-filtered dcpo and $\mathcal{K}$ a Scott compact saturated subset of $\mathcal{Q}(L)$. Then $\bigcup\mathcal{K}$ is compact saturated in $\Sigma L$.
Proof.
The saturation is clear. We only consider the compactness. Let $\mathcal{U}$ be a directed subset of $\sigma(L)$ with $\bigcup\mathcal{K}\subseteq\bigcup\mathcal{U}$. Then for each $K\in\mathcal{K}$, its compactness yields the existence of $U_{K}\in\mathcal{U}$ such that $K\subseteq U_{K}$, that is, $K\in\Box U_{K}$. Since the well-filteredness of $L$ implies that each $\Box U$ is Scott open in $\mathcal{Q}(L)$ and $\mathcal{U}$ is directed, $\{\Box U:U\in\mathcal{U}\}$ forms a directed family of Scott open subsets of $\mathcal{Q}(L)$ containing $\mathcal{K}$. So $\mathcal{K}\subseteq\Box U_{0}$ for some $U_{0}\in\mathcal{U}$ by its compactness. Thus $\bigcup\mathcal{K}\subseteq U_{0}$.
Theorem 3.12.
The endofunctor $\mathcal{Q}$ with the unit $\theta$ and the multiplication $\iota$ forms a monad on $\mathbf{WF}$, where $\theta$ and $\iota$ are defined as follows when they work on a well-filtered dcpo $L$,
$\theta_{L}(x)={\uparrow}x$, for any $x\in L$,
and
$\iota_{L}(\mathcal{K})=\bigcup\mathcal{K}$, for any $\mathcal{K}\in\mathcal{Q}(\mathcal{Q}(L))$.
Proof.
It is routine to check that $\mathcal{Q}$ is a monad.
It is noted that for a well-filtered dcpo $L$, arbitrary finite infs and directed sups in $\mathcal{Q}(L)$ are actually the finite unions and filtered intersections. Thus $\mathcal{Q}(L)$ is a meet continuous semilattice, in other words, $\mathcal{Q}(L)$ is a deflationary semilattice in $\mathbf{WF}$.
Lemma 3.13.
Every $\mathcal{Q}$-algebra $M$ is a deflationary semilattice in $\mathbf{WF}$.
Proof.
The definitions of $\mathcal{Q}$-algebras reveal that $M$ is a retract of $\mathcal{Q}(M)$. Then by [21, proposition 5.6], we know that $M$ is a deflationary semilattice in $\mathbf{WF}$.
Based on [21, Lemma 4.4] given by Andrea Schalk, we can derive the following results.
Lemma 3.14.
Let $(M,\alpha)$ be a $\mathcal{Q}$-algebra. Then for each $K\in\mathcal{Q}(M)$, its infimum $\wedge K$ exists.
Lemma 3.15.
Each $\mathcal{Q}$-algebra homomorphism preserves infs of compact sets, hence in particular finite infs.
Proposition 3.16.
The Eilenberg-Moore category of $\mathcal{Q}$ over $\mathbf{WF}$ is a subcategory of the category of deflationary semilattices in $\mathbf{WF}$.
Unfortunately, we cannot accurately characterize what the $\mathcal{Q}$-algebras are. Moreover, $\mathcal{Q}$ does not give the free deflationary semilattice for $\mathbf{WF}$ because there are deflationary semilattices in $\mathbf{WF}$ that are not $\mathcal{Q}$-algebras, as shown in the following example.
Example 3.17.
Let $\mathcal{L}$ be the dcpo given in Section 7.2.4 by Schalk in [21], that is,
$\mathcal{L}=\{A\subseteq[0,1]:A\neq\emptyset\ \mathrm{and}\ [0,1]\setminus A\ \mathrm{is\ countable}\}$.
It has been proved to be a deflationary semilattice in $\mathbf{DCPO}$. Besides, we can see that every nonempty finite subset of $\mathcal{L}$ has a sup, which is indeed the union of its elements. It then follows that $\mathcal{L}$ is an up-complete $\vee$-semilattice, this implies the well-filteredness of $\Sigma\mathcal{L}$ witnessed by [25, Proposition 3.1]. Thus $\mathcal{L}$ is a deflationary semilattice in $\mathbf{WF}$. However, there exists a compact saturated subset
$K={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\uparrow}}\{[0,1]\setminus\{x\}:x\in[0,1/2]\}$,
which has no infimum in $\mathcal{L}$.
Hence, by Lemma 3.14, $\mathcal{L}$ fails to be a $\mathcal{Q}$-algebra. So $\mathcal{Q}$ cannot give the free deflationary semilattice for $\mathbf{WF}$.
4 The commutativity between $\mathcal{H}$ and $\mathcal{Q}$
Definition 4.1.
Let $L$ be a well-filtered dcpo. We define $\phi_{L}:\mathcal{H}(\mathcal{Q}(L))\rightarrow\mathcal{Q}(\mathcal{H}(L))$ and $\psi_{L}:\mathcal{Q}(\mathcal{H}(L))\rightarrow\mathcal{H}(\mathcal{Q}(L))$ as
$\phi_{L}(\mathcal{A})=\{A\in\mathcal{H}(L):\forall K\in\mathcal{A},A\cap K\neq\emptyset\}$, for any $\mathcal{A}\in\mathcal{H}(\mathcal{Q}(L))$,
$\psi_{L}(\mathcal{K})=\{K\in\mathcal{Q}(L):\forall A\in\mathcal{K},K\cap A\neq\emptyset\}$, for any $\mathcal{K}\in\mathcal{Q}(\mathcal{H}(L))$.
Notably, the forms of these two definitions are similar to those in [4] and [11], where the objects they discussed are quasi-polish spaces and $\mathcal{U}_{S}$-admitting dcpos, respectively. For convenience, we will omit the subscripts when the object acted on is clear.
Lemma 4.2.
Every complete lattice endowed with the Scott topology is coherent.
Proof.
[25, Corollary 3.2] guarantees the well-filteredness of a complete lattice $L$. Since for any $x,y\in L$, ${\uparrow}x\cap{\uparrow}y={\uparrow}(x\vee y)$, by [15, Corollary 3.2], we have that $L$ is coherent.
Lemma 4.3.
The functions $\phi$ and $\psi$ defined in Definition 4.1 are well-defined.
Proof.
To make it clear, let us define $\phi^{\prime}:\mathcal{Q}(L)\rightarrow\mathcal{Q}(\mathcal{H}(L))$ and $\psi^{\prime}:\mathcal{H}(L)\rightarrow\mathcal{H}(\mathcal{Q}(L))$ as follows:
$\phi^{\prime}(K)=\{A\in\mathcal{H}(L):A\cap K\neq\emptyset\}$, for any $K\in\mathcal{Q}(L)$,
$\psi^{\prime}(A)=\{K\in\mathcal{Q}(L):K\cap A\neq\emptyset\}$, for any $A\in\mathcal{H}(L)$.
Then $\phi(\mathcal{A})=\bigcap\{\phi^{\prime}(K):K\in\mathcal{A}\}$ and $\psi(\mathcal{K})=\bigcap\{\psi^{\prime}(A):A\in\mathcal{K}\}$. One can easily check that $\phi^{\prime}(K)=\mathcal{Q}(\eta_{L})(K)={\uparrow}\eta_{L}(K)$ and $\psi^{\prime}(A)=\mathcal{H}(\theta_{L})(A)=cl(\theta_{L}(A))$, which implies that $\phi^{\prime}$ and $\psi^{\prime}$ are well-defined. We know the intersection of arbitrary closed subsets is still closed. So, $\psi$ is well-defined. Now we focus on $\phi$. Lemma 4.2 reveals that $\Sigma\mathcal{H}(L)$ is coherent. Combining with the well-filteredness of $\Sigma\mathcal{H}(L)$, it will be a fact that each intersection of a collection of compact saturated subsets in $\mathcal{H}(L)$ is still compact saturated. Thus $\phi(\mathcal{A})\in\mathcal{Q}(\mathcal{H}(L))$, i.e., $\phi$ is well-defined.
Proposition 4.4.
Let $L$ be a well-filtered dcpo. The function $\phi:\mathcal{HQ}(L)\rightarrow\mathcal{QH}(L)$ defined in Definition 4.1 is Scott-continuous.
Proof.
Assume that there are $\mathcal{A},\mathcal{B}\in\mathcal{HQ}(L)$ with $\mathcal{A}\leq\mathcal{B}$, i.e., $\mathcal{A}\subseteq\mathcal{B}$. By the definition of $\phi$, one can easily see that $\phi(\mathcal{B})\subseteq\phi(\mathcal{A})$, equivalently, $\phi(\mathcal{A})\leq\phi(\mathcal{B})$. So, $\phi$ is order preserving. Let $\{\mathcal{A}_{i}:i\in I\}$ be a directed subset of $\mathcal{HQ}(L)$. To complete this proof, we need to verify that $\phi(\bigvee_{i\in I}\mathcal{A}_{i})=\bigvee_{i\in I}\phi(\mathcal{A}_{i})$, i.e., $\phi(cl(\bigcup_{i\in I}\mathcal{A}_{i}))=\bigcap_{i\in I}\phi(\mathcal{A}_{i})$. Since $\phi$ is order preserving, $\bigcap_{i\in I}\phi(\mathcal{A}_{i})\subseteq\phi(cl(\bigcup_{i\in I}\mathcal{A}_{i}))$ clearly holds. Now pick $A\in\phi(cl(\bigcup_{i\in I}\mathcal{A}_{i}))$, then for each $K\in\bigcup_{i\in I}\mathcal{A}_{i},A\cap K\neq\emptyset$. So we have $A\cap K\neq\emptyset,\forall i\in I,\forall K\in\mathcal{A}_{i}$, which immediately implies that $A\in\phi(\mathcal{A}_{i}),\forall i\in I$. Therefore, $A\in\bigcap_{i\in I}\phi(\mathcal{A}_{i})$ consequently, $\phi(cl(\bigcup_{i\in I}\mathcal{A}_{i}))\subseteq\bigcup_{i\in I}\phi(\mathcal{A}_{i})$ holds.
Definition 4.5.
A poset $P$ is said to have property $(\mathrm{KC})$ if for every compact saturated subset $\mathcal{K}$ of $\Gamma(P)$ under the Scott topology, any $U\in\sigma(P)$ that meets all members of $\mathcal{K}$ contains a compact saturated subset $K$ that still meets all members of $\mathcal{K}$.
Recall that a space $X$ is consonant if and only if for any Scott open subset $\mathcal{U}\in\mathcal{O}(X)$ and any $U\in\mathcal{U}$, there exists a $K\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{Q}(X)}$ such that $U\in\Phi(K)\subseteq\mathcal{U}$, where $\Phi(K)=\{V\in\mathcal{O}(X):K\subseteq V\}$. The concept was proposed initially by Dolecki et al. in [5] to answer some questions related to topologies on the hyperspace of closed subsets of a topological space. We say a poset is consonant if it endowed with the Scott topology is consonant.
Lemma 4.6.
If a poset $P$ is consonant, then it satisfies the property $(\mathrm{KC})$.
Proof.
Let $\mathcal{K}$ be a compact saturated subset of $(\Gamma(P),\sigma(\Gamma(P)))$ and $U$ a Scott open subset of $P$ satisfying that $U\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. Then for each $A\in\mathcal{K}$, $A\in\Diamond U$, which implies that $\mathcal{K}\subseteq\Diamond U$, i.e., $\mathcal{K}\in\Box\Diamond U$. Notably, [4, Theorem 6.13] tells us that the consonance of a topological space $X$ implies the equation $\tau^{-1}(\Diamond\Box U)=\Box\Diamond U$ holds. For the space $\Sigma P$, $\tau$ in [4, Theorem 6.13] is just $\psi$ defined by us. So the consonance of $\Sigma P$ will indicate that $\Box\Diamond U\subseteq\psi^{-1}(\Diamond\Box U)$. Thus $\mathcal{K}\in\psi^{-1}(\Diamond\Box U)$. From which we have that $\psi(\mathcal{K})\in\Diamond\Box U$, i.e., $\psi(\mathcal{K})\cap\Box U\neq\emptyset$. So there exists a $K\in\psi(\mathcal{K})$ contained in $U$. Therefore, we find a compact saturated subset $K$ contained in $U$ that meets each element of $\mathcal{K}$. Hence, $P$ has the property $(\mathrm{KC})$.
Now see [4, Theorem 6.10] again, it states that for a space $X$, $\sigma(\mathcal{O}(X))$ and $\mathcal{Q}(\Gamma(X))$ are isomorphic lattices. Recently, Chen et al. in [3] proved that if a poset $P$ under the Scott topology is core-compact, then $\sigma(\Gamma(P))=\nu(\Gamma(P))$. Thus for a poset $P$, $\sigma(\sigma(P))$ would be isomorphic to $\mathcal{Q}(\Gamma(P))$ as long as $\Sigma P$ is core-compact. Combining all of these results and the above lemma, we obtain the following result.
Corollary 4.7.
Let $P$ be a poset. If $(P,\sigma(P))$ is core-compact and has the property $(\mathrm{KC})$, then $P$ is consonant.
Theorem 4.8.
Let $L$ be a well-filtered dcpo. Then $\mathcal{H}(\mathcal{Q}(L))$ is isomorphic to $\mathcal{Q}(\mathcal{H}(L))$ under the maps $\phi$ and $\psi$ if and only if the following conditions are satisfied.
  1. 1.
    $\sigma(Q(L))\subseteq\upsilon(Q(L));$ and
  2. 2.
    $L$ has the property $(\mathrm{KC})$.
Proof.
$(\Rightarrow)$: Let $\mathcal{U}$ be a Scott open subset of $\mathcal{Q}(L)$ and $K$ an element in $\mathcal{U}$. Then $K$ does not belong to the Scott closed subset $\mathcal{Q}(L)\setminus\mathcal{U}$. Since $\psi$ is surjective, there exists a $\mathcal{K}\in\mathcal{Q}(\mathcal{H}(L))$ such that $\mathcal{Q}(L)\setminus\mathcal{U}=\psi(\mathcal{K})$, then $K\notin\psi(\mathcal{K})$. By the definition of $\psi$, it yields that $K\cap A=\emptyset$ for some $A\in\mathcal{K}$. The Scott open subset $V=L\setminus A$ contains $K$, that is, $K\in\Box V$. Thus we only need to show $\Box V\subseteq\mathcal{U}$. For any $M\in\Box V$, $M\cap A=\emptyset$, which implies that $M\notin\psi(\mathcal{K})$. Hence, $M\in\mathcal{U}$.
Let $\mathcal{K}$ be a Scott compact saturated subset of $\Gamma(L)$ and $U\in\sigma(L)$ satisfying $U\cap A\neq\emptyset$ for all $A\in\mathcal{K}$. Because of $U\cap(L\setminus U)=\emptyset$, we have $L\setminus U\notin\mathcal{K}$. The fact that $\phi$ is surjective suggests that there is an $\mathcal{A}\in\mathcal{H}(\mathcal{Q}(L))$ such that $\mathcal{K}=\phi(\mathcal{A})$. Thus $L\setminus U\notin\phi(\mathcal{A})$. This means $(L\setminus U)\cap K_{0}=\emptyset$ i.e., $K_{0}\subseteq U$ for some $K_{0}\in\mathcal{A}$. Now suppose that $K_{0}\cap A_{1}=\emptyset$ for some $A_{1}\in\mathcal{K}$. Then $A_{1}\notin\phi(\mathcal{A})=\mathcal{K}$, which is an contradiction obviously. Thus $K_{0}$ intersects each $A\in\mathcal{K}$ and $L$ has the property $(\mathrm{KC})$ as desired.
$(\Leftarrow)$: Obviously, $\phi$ and $\psi$ are order-preserving. Since order-preserving isomorphisms between dcpos are Scott-continuous, the proof will be complete after showing that $\psi\circ\phi$ and $\phi\circ\psi$ are identities, more precisely, the following two equations hold (for arbitrary $\mathcal{A}\in\mathcal{H}(\mathcal{Q}(L))$ and $\mathcal{K}\in\mathcal{Q}(\mathcal{H}(L))$):
$\psi\circ\phi(\mathcal{A})=\{K\in\mathcal{Q}(L):\forall A\in\phi(\mathcal{A}),K\cap A\neq\emptyset\}=\mathcal{A}$,
$\phi\circ\psi(\mathcal{K})=\{A\in\mathcal{H}(L):\forall K\in\psi(\mathcal{K}),A\cap K\neq\emptyset\}=\mathcal{K}$.
Note that the definitions of $\phi$ and $\psi$ determine that $\mathcal{A}\subseteq\psi\circ\phi(\mathcal{A})$ and $\mathcal{K}\subseteq\phi\circ\psi(\mathcal{K})$. Thus we just need to check the reverse.
As to the first one, suppose that there is a $K\in\mathcal{Q}(L)$ belonging to $\psi\circ\phi(\mathcal{A})$ but not to $\mathcal{A}$. Then $K\in\mathcal{Q}(L)\setminus\mathcal{A}$, which is Scott open in $\mathcal{Q}(L)$ and further an open subset of $\mathcal{Q}(L)$ with respect to the upper Vietoris topology by $(\mathrm{i})$. Thus there exists a $U\in\sigma(L)$ such that $K\in\Box U\subseteq\mathcal{Q}(L)\setminus\mathcal{A}$, which implies $K\cap(L\setminus U)=\emptyset$. This will contradict the fact that $K\in\psi\circ\phi(\mathcal{A})$ if we could show $L\setminus U\in\phi(\mathcal{A})$. Assume $(L\setminus U)\cap K_{0}=\emptyset$ for some $K_{0}\in\mathcal{A}$. Then $K_{0}\in\Box U$, thus $K_{0}\notin\mathcal{A}$, a contradiction. So $L\setminus U\in\phi(\mathcal{A})$. Therefore, $\psi\circ\phi(\mathcal{A})\subseteq\mathcal{A}$.
Regarding the second, for the sake of a contradiction, we assume that there is an $A\in\phi\circ\psi(\mathcal{K})$ but $A\notin\mathcal{K}$. It suggests that each $C\in\mathcal{K}$ is not contained in $A$, that is to say, $(L\setminus A)\cap C\neq\emptyset$ for all $C\in\mathcal{K}$. Since $L$ has the property $(\mathrm{KC})$, we can find a $K_{0}\subseteq L\setminus A$ intersecting each $C$ in $\mathcal{K}$. Then $K_{0}\in\psi(\mathcal{K})$ and $K_{0}\cap A=\emptyset$, which contradicts that $A\in\phi\circ\psi(\mathcal{K})$. Thus $A\in\mathcal{K}$, which obviously leads to a contradiction. So our assumption is wrong and $\phi\circ\psi(\mathcal{K})\subseteq\mathcal{K}$ holds.
It was Schalk who proved that for a locally compact sober space $X$, the upper Vietoris topology and the Scott topology on $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{Q}(X)$} coincide (see [21, Lemma 7.26]), meanwhile, [3, Proposition 5.4] reveals that each locally compact space is consonant. Therefore, combining Lemma 4.6 and Theorem 4.8, we further draw the conclusion in the following.
Corollary 4.9.
If $L$ is a locally compact sober dcpo, then $\mathcal{QH}(L)$ is isomorphic to $\mathcal{HQ}(L)$ under the function $\phi$ (or $\psi$).
5 The double power well-filtered dcpo
During the study of monads, there has always been a problem worth thinking about: Given two monads with underlying functors $S$ and $T$ on a category $\mathbf{C}$, does $T\circ S$ (or $S\circ T$) still carry the structure of a monad? Until 1969, Beck introduced the concepts of distributive laws and showed that the existence of distributive laws provides a sufficient condition for such a composition to form a monad [2].
Definition 5.1.
([2]) Let $\mathcal{S}=(\mathcal{S},\eta^{\mathcal{S}},\mu^{\mathcal{S}})$ and $\mathcal{T}=(\mathcal{T},\eta^{\mathcal{T}},\mu^{\mathcal{T}})$ be two monads on the category $\mathbf{C}$. A distributive law of $\mathcal{S}$ over $\mathcal{T}$ is a natural transformation $l:\mathcal{TS}\rightarrow\mathcal{ST}$ such that the following diagrams commute (Figure 1).
[…]
Figure 1
Definition 5.2.
([1]) Let $\mathcal{T}=(\mathcal{T},\eta,\mu)$ and $\mathcal{T}^{\prime}=(\mathcal{T}^{\prime},\eta^{\prime},\mu^{\prime})$ be two monads. A morphism of monads is a natural transformation $\lambda:\mathcal{T}\rightarrow\mathcal{T}^{\prime}$ such that the following two diagrams commute (Figure 2).
[…]
Figure 2
Particularly, the morphisms of monads are called triple maps by Beck in [2]. Meanwhile, Beck also proved the following theorem.
Theorem 5.3.
([2]) Let $\mathcal{S}=(\mathcal{S},\eta^{\mathcal{S}},\mu^{\mathcal{S}})$, $\mathcal{T}=(\mathcal{T},\eta^{\mathcal{T}},\mu^{\mathcal{T}})$ be two monads on the category $\mathbf{C}$. Then the following statements are equivalent:
  1. 1.
    There are distributive laws $d:\mathcal{ST}\rightarrow\mathcal{TS};$
  2. 2.
    There are multiplications $\mu:\mathcal{TSTS}\rightarrow\mathcal{TS}$ such that
    • (a)
      $(\mathcal{TS},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\eta_{\mathcal{S}}^{\mathcal{T}}\eta^{\mathcal{S}}},\mu)$ is a monad;
    • (b)
      The natural transformations $\eta^{\mathcal{T}}_{\mathcal{S}}:\mathcal{S}\rightarrow\mathcal{TS}$ and $\mathcal{T}\eta^{\mathcal{S}}:\mathcal{T}\rightarrow\mathcal{TS}$ are morphisms of monads.
  3. 3.
    Liftings of $\tilde{\mathcal{S}}$ of the monad $\mathcal{S}$ into $\mathbf{C}^{\mathcal{T}}$.
From the proposition presented in Section 2 of Beck [2], a more general result concerning the morphisms of monads was established by Fritz et al.
Lemma 5.4.
([7]) Let $\mathcal{S}=(\mathcal{S},\eta_{\mathcal{S}},\mu_{\mathcal{S}})$ and $\mathcal{T}=(\mathcal{T},\eta_{\mathcal{T}},\mu_{\mathcal{T}})$ be two monads on a category $\mathbf{C}$ and $\lambda:\mathcal{S}\rightarrow\mathcal{T}$ a morphism of monads. Then every $\mathcal{T}$-algebra $(A,\alpha)$ can be equipped with an $\mathcal{S}$-algebra via $(A,\alpha)\mapsto(A,\alpha\circ\lambda)$. Moreover, a $\mathcal{T}$-algebra homomorphism $f:(A,\alpha)\rightarrow(B,\beta)$ in duces an $\mathcal{S}$-algebra homomorphism $f:(A,\alpha\circ\lambda)\rightarrow(B,\beta\circ\lambda)$ in a functorial way.
Proposition 5.5.
$\phi$ : $\mathcal{H}\mathcal{Q}\rightarrow\mathcal{Q}\mathcal{H}$ defined in Definition 4.1 is a distributive law of $\mathcal{Q}$ over $\mathcal{H}$.
Proof.
$\bf{Claim~1:}$ $\phi$ is a natural transformation.
Assume $L$ and $M$ are well-filtered dcpos and $f:L\rightarrow M$ is a Scott-continuous mapping. We need to show the following diagram commutes,
$\mathcal{HQ}(L)$$\mathcal{QH}(L)$$\phi_{L}$$\mathcal{HQ}(M)$$\mathcal{QH}(M)$$\mathcal{HQ}(f)$$\mathcal{QH}(f)$$\phi_{M}$
i.e., for any $\mathcal{A}\in\mathcal{HQ}(L)$, $\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})=\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$, where
$\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})={\uparrow}\{cl(f(A)):\forall K\in\mathcal{A},A\cap K\neq\emptyset\}$,
$\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})=\{B\in\mathcal{H}(M):\forall N\in cl\{{\uparrow}f(K):K\in\mathcal{A}\},B\cap N\neq\emptyset\}$.
First, suppose that there exists a $cl(f(A))\in\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})$ but $cl(f(A))\notin\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$, that is, we can find a $N_{0}\in cl\{{\uparrow}f(K):K\in\mathcal{A}\}$ satisfying $cl(f(A))\cap N_{0}=\emptyset$. From which we can see that $N_{0}\in\Box(M\setminus cl(f(A)))$. Since the well-filteredness of $M$ implies that $\Box(M\setminus cl(f(A)))\in\sigma(\mathcal{Q}(M))$, $\Box(M\setminus cl(f(A)))\cap\{{\uparrow}f(K):K\in\mathcal{A}\}\neq\emptyset$. So there is a $K_{0}\in\mathcal{A}$ such that ${\uparrow}f(K_{0})\subseteq M\setminus cl(f(A))$, that is to say, ${\uparrow}f(K_{0})\cap cl(f(A))=\emptyset$. Then $K_{0}\cap A=\emptyset$, which obviously contradicts the fact that $A$ meets every element in $\mathcal{A}$. Thus, $\{cl(f(A)):\forall K\in\mathcal{A},A\cap K\neq\emptyset\}\subseteq\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$; furthermore, $\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})\subseteq\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$.
Now we prove the reverse inclusion. Let $B\in\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$. Then for each $K\in\mathcal{A}$, $B\cap{\uparrow}f(K)\neq\emptyset$, combing with the downness of $B$, which further indicates that $B\cap f(K)\neq\emptyset$. We pick $x_{K}\in K$ such that $f(x_{K})\in B\cap f(K)$ and set $A_{0}=cl(\{x_{K}:K\in\mathcal{A}\})$. Then $cl(f(A_{0}))=cl(f(\{x_{K}:K\in\mathcal{A}\}))=cl(\{f(x_{K}):K\in\mathcal{A}\})\subseteq B$ by the closeness of $B$. Meanwhile, $A_{0}\cap K\neq\emptyset$ for each $K\in\mathcal{A}$. So $B\in\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})$, then, $\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})\subseteq\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})$ holds.
$\bf{Claim~2:}$ The first diagram in Definition 5.1 commutes.
Concretely, given a well-filtered dcpo $L$, we need to demonstrate that the following diagram commutes.
$\mathcal{H}(L)$$\mathcal{HQ}(L)$$\mathcal{H}\theta_{L}$$\mathcal{QH}(L)$$\theta_{\mathcal{H}(L)}$$\phi_{L}$
Let $A\in\mathcal{H}(L)$. Then $\theta_{\mathcal{H}(L)}(A)={\uparrow}\{A\}$ and
$\phi_{L}\circ\mathcal{H}\theta_{L}(A)=\{B\in\mathcal{H}(L):\forall K\in cl(\{{\uparrow}x:x\in A\}):B\cap K\neq\emptyset\}$.
First, suppose that there is a $K\in cl(\{{\uparrow}x:x\in A\})$ such that $A\cap K=\emptyset$, then $K\subseteq L\setminus A$, i.e., $K\in\Box(L\setminus A)$. By the well-filteredness of $L$, we know $\Box(L\setminus A)\in\sigma(\mathcal{Q}(L))$. So $\Box(L\setminus A)\cap\{{\uparrow}x:x\in A\}\neq\emptyset$, that is to say, we can find an $x\in A$ such that ${\uparrow}x\subseteq L\setminus A$. Thus $x\notin A$, clearly, which is a contradiction. So for any $K\in cl(\{{\uparrow}x:x\in A\}),A\cap K\neq\emptyset$, which implies that $A\in\phi_{L}\circ\mathcal{H}\theta_{L}(A)$; hence, $\theta_{\mathcal{H}(L)}(A)={\uparrow}\{A\}\subseteq\phi_{L}\circ\mathcal{H}\theta_{L}(A)$.
Now consider the reverse inclusion. For any $B\in\phi_{L}\circ\mathcal{H}\theta_{L}(A)$, we have $B\cap{\uparrow}x\neq\emptyset,\forall x\in A$. It means that each $x\in A$ belongs to $B$. Thus $A\subseteq B$, equivalently, $B\in{\uparrow}\{A\}$. By the arbitrariness of $B$, we conclude that $\phi_{L}\circ\mathcal{H}\theta_{L}(A)\subseteq\theta_{\mathcal{H}(L)}(A)$.
$\bf{Claim~3:}$ The second diagram in Definition 5.1 commutes.
Specifically, given a well-filtered dcpo $L$, we need to prove that the following diagram commutes,
$\mathcal{Q}(L)$$\mathcal{HQ}(L)$$\eta_{\mathcal{Q}(L)}$$\mathcal{QH}(L)$$\mathcal{Q}\eta_{L}$$\phi_{L}$
i.e., to verify $\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)=\mathcal{Q}\eta_{L}(K)$ for any $K\in\mathcal{Q}(L)$, where $\mathcal{Q}\eta_{L}(K)={\uparrow}\{{\downarrow}x:x\in K\}$ and
$\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)=\{A\in\mathcal{H}(L):\forall M\in{\downarrow}\{K\}:A\cap M\neq\emptyset\}$.
Let $A$ be an arbitrary element in $\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)$. Then $A\cap K\neq\emptyset$, and thus, there is an $x\in A\cap K$. By the closedness of $A$, we have ${\downarrow}x\subseteq A$, i.e., $A\in{\uparrow}\{{\downarrow}x\}$. So $A\in\mathcal{Q}\eta_{L}(K)$. The arbitrariness of $A$ guarantees that $\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)\subseteq\mathcal{Q}\eta_{L}(K)$.
Now focus on each $x\in K$, we have ${\downarrow}x\cap K\neq\emptyset$. So for any $M\in{\downarrow}\{K\}$, i.e., $K\subseteq M$, ${\downarrow}x\cap M$ always holds. This means $\{{\downarrow}x:x\in K\}\subseteq\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)$; thus, $\mathcal{Q}\eta_{L}(K)\subseteq\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)$.
$\bf{Claim~4:}$ The third diagram in Definition 5.1 commutes, i.e., the following one commutes:
$\mathcal{HQQ}(L)$$\mathcal{QHQ}(L)$$\mathcal{QQH}(L)$$\phi_{\mathcal{Q}(L)}$$\mathcal{Q}\phi_{L}$$\mathcal{HQ}(L)$$\mathcal{QH}(L)$$\mathcal{H}\tau_{L}$$\tau_{\mathcal{H}(L)}$$\phi_{L}$
For any $\mathbf{A}\in\mathcal{HQQ}(L)$, by computation, we have
$\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})=\bigcup{\uparrow}\{\phi_{L}(\mathcal{A}):\forall\mathcal{K}\in\mathbf{A},\mathcal{A}\cap\mathcal{K}\neq\emptyset\}$,
$\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})=\{A\in\mathcal{H}(L):\forall K\in cl(\{\bigcup\mathcal{K}:\mathcal{K}\in\mathbf{A}\}),A\cap K\neq\emptyset\}$.
Consider $\phi_{L}(\mathcal{A})=\{B\in\mathcal{H}(L):\forall M\in\mathcal{A},B\cap M\neq\emptyset\}$, where $\mathcal{A}$ meets each element of $\mathbf{A}$. Suppose that there exists a $B\in\mathcal{H}(L)$ that belongs to $\phi_{L}(\mathcal{A})$ but not to $\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})$, i.e., there is a $K\in cl(\{\bigcup\mathcal{K}:\mathcal{K}\in\mathbf{A}\})$ satisfying that $B\cap K=\emptyset$. It further implies that $K\in\Box(L\setminus B)$. Applying the well-filteredness of $L$ again, we obtain that $\Box(L\setminus B)$ is a Scott open subset of $\mathcal{Q}(L)$; thus, $\Box(L\setminus B)\cap\{\bigcup\mathcal{K}:\mathcal{K}\in\mathbf{A}\}\neq\emptyset$. So we can find a $\mathcal{K}\in\mathbf{A}$ satisfying $\bigcup\mathcal{K}\subseteq L\setminus B$. From which we say that for any $C\in\mathcal{K}$, $C\cap B=\emptyset$. Combining with the fact that $B\in\phi_{L}(\mathcal{A})$, one can conclude that every $C\in\mathcal{K}$ is not in $\mathcal{A}$, that is to say, $\mathcal{K}\cap\mathcal{A}=\emptyset$. However, $\mathcal{A}$ meets every element of $\mathbf{A}$, a contradiction. Thus, each $\phi_{L}(\mathcal{A})$ is contained in $\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})$; accordingly, $\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})\subseteq\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})$.
Take an arbitrary element $A\in\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})$, then for each $\mathcal{K}\in\mathbf{A},A\cap\bigcup\mathcal{K}\neq\emptyset$. So there is an $a_{\mathcal{K}}\in A\cap\bigcup\mathcal{K}$, which further indicates that ${\uparrow}a_{\mathcal{K}}\subseteq K$, i.e., $K\leq{\uparrow}a_{\mathcal{K}}$ for some $K\in\mathcal{K}$. Since $\mathcal{K}$ is an upper set, ${\uparrow}a_{\mathcal{K}}\in\mathcal{K},\forall\mathcal{K}\in\mathbf{A}$. We set $C=cl\{a_{\mathcal{K}}:\mathcal{K}\in\mathbf{A}\}$. Then under the mapping $\mathcal{H}\theta_{L}$, we can obtain a Scott closed subset $cl(\{{\uparrow}c:c\in C\})$ of $\mathcal{Q}(L)$. Next, we verify that $A\in\phi_{L}(cl(\{{\uparrow}c:c\in C\}))$ and $\phi_{L}(cl(\{{\uparrow}c:c\in C\}))\in\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})$. First, if there is an $M\in cl(\{{\uparrow}c:c\in C\})$ such that $A\cap M=\emptyset$, then $M\in\Box(L\setminus A)$. Since the well-filteredness of $L$ implies the Scott openness of $\Box(L\setminus A)$, $\Box(L\setminus A)\cap\{{\uparrow}c:c\in C\}\neq\emptyset$. That is, there exists a $c\in C$ such that ${\uparrow}c\subseteq L\setminus A$, then, $c\in L\setminus A$. As $L\setminus A\in\sigma(L)$ and $c\in cl\{a_{\mathcal{K}}:\mathcal{K}\in\mathbf{A}\}$, we can find an $a_{\mathcal{K}}$ such that $a_{\mathcal{K}}\in L\setminus A$. i.e., $a_{\mathcal{K}}\notin A$, obviously, this leads to a contradiction. Thus, $A\in\phi_{L}(cl(\{{\uparrow}c:c\in C\}))$. Then, since for each $a_{\mathcal{K}}\in C$, ${\uparrow}a_{\mathcal{K}}\in cl(\{{\uparrow}c:c\in C\})$ and ${\uparrow}a_{\mathcal{K}}\in\mathcal{K}$, $cl(\{{\uparrow}c:c\in C\})\cap\mathcal{K}\neq\emptyset,\forall\mathcal{K}\in\mathbf{A}$. So $\phi_{L}(cl(\{{\uparrow}c:c\in C\}))\in\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})$. As a conclusion, we find an element $\phi_{L}(cl(\{{\uparrow}c:c\in C\}))$ in $\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})$ such that $A$ belongs to it, by the arbitrariness of $A$, we get that $\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})\subseteq\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})$ holds.
$\bf{Claim~5:}$ The forth diagram in Definition 5.1 commutes, i.e., the following one commutes:
$\mathcal{HHQ}(L)$$\mathcal{HQH}(L)$$\mathcal{QHH}(L)$$\mathcal{H}\phi_{L}$$\phi_{\mathcal{H}(L)}$$\mathcal{HQ}(L)$$\mathcal{QH}(L)$$\mu_{\mathcal{Q}(L)}$$\mathcal{Q}\mu_{L}$$\phi_{L}$
Let $\mathbf{A}$ be an arbitrary element in $\mathcal{HHQ}(L)$. By computation, we have
$\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})=\{A\in\mathcal{H}(L):\forall K\in\bigcup\mathbf{A},A\cap K\neq\emptyset\}$,
$\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})={\uparrow}\{\bigcup\mathcal{C}:\mathcal{C}\in\mathcal{HH}(L),\forall\mathcal{K}\in cl(\{\phi_{L}(\mathcal{A}):\mathcal{A}\in\mathbf{A}\}),\mathcal{C}\cap\mathcal{K}\neq\emptyset\}$.
For each $A\in\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})$, we have ${\downarrow}\{A\}\in\mathcal{HH}(L)$ and $A=\bigcup{\downarrow}\{A\}$. Suppose that there exists a $\mathcal{K}\in cl(\{\phi_{L}(\mathcal{A}):\mathcal{A}\in\mathbf{A}\})$ such that ${\downarrow}\{A\}\cap\mathcal{K}=\emptyset$, then $\mathcal{K}\in\Box(\mathcal{H}(L)\setminus{\downarrow}\{A\})$. By the well-filteredness of $\mathcal{H}(L)$, we know $\Box(\mathcal{H}(L)\setminus{\downarrow}\{A\})$ is a Scoot open subset of $\mathcal{QH}(L)$. Therefore, $\Box(\mathcal{H}(L)\setminus{\downarrow}\{A\})\cap\{\phi_{L}(\mathcal{A}):\mathcal{A}\in\mathbf{A}\}\neq\emptyset$. So there is an $\mathcal{A}\in\mathbf{A}$ such that $\phi_{L}(\mathcal{A})\subseteq\mathcal{H}(L)\setminus{\downarrow}\{A\}$. This further indicates that $\phi_{L}(\mathcal{A})\cap{\downarrow}\{A\}=\emptyset$, i.e., $A\notin\phi_{L}(\mathcal{A})$. From which we derive that $A\cap K=\emptyset$ for some $K\in\mathcal{A}$. Note that $\mathcal{A}\in\mathbf{A}$ implies $K\in\bigcup\mathbf{A}$. As a result, clearly, $A\cap K=\emptyset$ contradicts the fact that $A\in\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})$. So $A=\bigcup{\downarrow}\{A\}\in\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})$, and hence, $\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})\subseteq\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})$.
Pick an arbitrary element $\bigcup\mathcal{C}$, where $\mathcal{C}\in\mathcal{HH}(L)$ and for each $\mathcal{K}\in cl(\{\phi_{L}(\mathcal{A}):\mathcal{A}\in\mathbf{A}\}),\mathcal{C}\cap\mathcal{K}\neq\emptyset$. Assume that there exists a $K\in\bigcup\mathbf{A}$, i.e., $K\in\mathcal{A}$ for some $\mathcal{A}\in\mathbf{A}$, such that $\bigcup\mathcal{C}\cap K=\emptyset$. Then for each $C\in\mathcal{C}$, $C\cap K=\emptyset$, which immediately means that $C\notin\phi_{L}(\mathcal{A}),\forall C\in\mathcal{C}$, in other words, $\mathcal{C}\cap\phi_{L}(\mathcal{A})=\emptyset$. This contradicts $\bigcup\mathcal{C}\in\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})$. So $\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})\subseteq\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})$.
To sum up, $\phi$ defined in Definition 4.1 is a distributive law of $\mathcal{Q}$ over $\mathcal{H}$.
From the above results, we derive that $\mathcal{Q}\mathcal{H}$ is a monad on $\mathbf{WF}$. Notably, Steven Vickers studied the Eilenberg-Moore algebras of the double powerlocale $P_{U}P_{L}$ (where $P_{U}$ and $P_{L}$ denote the upper and lower powerlocales, which commute up to isomorphisms via a distributive law) and it is concluded that $P_{U}P_{L}$-algebras correspond to localic frames [24]. Motivated by this work, we now investigate the algebraic structures of $\mathcal{Q}\mathcal{H}$-algebras.
We first give the unit $\gamma$ and the multiplication $\rho$ of $\mathcal{Q}\mathcal{H}$ with the following diagrams, which are closely related to that of $\mathcal{Q}$ and $\mathcal{H}$ according to Beck [2].
[…]
Definition 5.6.
([8]) Let $\mathbf{C}$ be a poset-enriched category. A monad $\mathcal{T}=(\mathcal{T},\eta,\mu)$ on $\mathbf{C}$ is called a right $\rm{KZ}$-monad if $\mathcal{T}$ is a poset-functor and satisfies $\eta_{TX}\leq T\eta_{X}$ for all $X\in\mathbf{C}$. Left $\rm{KZ}$-monads are defined poset-dually, by reversing the inequality.
Remark 5.7.
[…]
  1. 1.
    For all $C\in\mathbf{C}$, an arrow $\alpha:\mathcal{T}C\rightarrow C$ is a structure map iff $\eta_{C}\dashv\alpha$ and $\alpha\circ\eta_{C}=id_{C}$.
  2. 2.
    $\eta_{\mathcal{T}C}\dashv\mu_{C}$ for all $C\in\mathbf{C}$.
  3. 3.
    $\mu_{C}\dashv\mathcal{T}\eta_{C}$ for all $C\in\mathbf{C}$.
Proposition 5.8.
[8] Let $\mathcal{T}=(\mathcal{T},\eta,\mu)$ be a $\rm{KZ}$-monad on the category $\mathbf{C}$. Then by $(\mathrm{KZ}_{1})$, every object has at most one structure map.
Lemma 5.9.
The category $\mathbf{WF}$ is a poset-enriched category; $\mathcal{H}$ and $\mathcal{Q}$ are both poset-functors.
Proposition 5.10.
$\mathcal{H}$ is a left $\rm{KZ}$-monad and $\mathcal{Q}$ is a right $\rm{KZ}$-monad.
Proof.
The definitions of the two $\rm{KZ}$-monads make the proof trivial.
Proposition 5.11.
Let $(L,\alpha)$ be a $\mathcal{Q}\mathcal{H}$-algebra in $\mathbf{WF}$. Then
  1. 1.
    $(\mathcal{Q}\mathcal{H}(L),\iota_{\mathcal{H}(L)})$ and $(L,\alpha^{\mathcal{Q}})$ with $\alpha^{\mathcal{Q}}=\alpha\circ\mathcal{Q}\eta_{L}$ are $\mathcal{Q}$-algebras, and $\alpha:(\mathcal{Q}\mathcal{H}(L),\iota_{\mathcal{H}(L)})\rightarrow(L,\alpha^{\mathcal{Q}})$ is a $\mathcal{Q}$-algebra homomorphism.
  2. 2.
    $(\mathcal{Q}\mathcal{H}(L),\rho_{L}\circ\theta_{\mathcal{H}(L)})$ and $(L,\alpha^{\mathcal{H}})$ with $\alpha^{\mathcal{H}}=\alpha\circ\theta_{\mathcal{H}(L)}$ are $\mathcal{H}$-algebras and $\alpha:(\mathcal{Q}\mathcal{H}(L),\rho_{L}\circ\theta_{\mathcal{H}(L)})\rightarrow(L,\alpha^{\mathcal{H}})$ is an $\mathcal{H}$-algebra homomorphism.
  3. 3.
    $\alpha$ is the unique $\mathcal{Q}\mathcal{H}$-structure map of $L$.
Proof.
By Proposition 2.4 (1) and (2), $(\mathcal{Q}\mathcal{H}(L),\rho_{L})$ is a $\mathcal{Q}\mathcal{H}$-algebra and $\alpha:(\mathcal{Q}\mathcal{H}(L),\rho_{L})\rightarrow(L,\alpha)$ is a $\mathcal{Q}\mathcal{H}$-algebra homomorphism.
  1. 1.
    Theorem 5.3 and Proposition 5.5 imply that $\mathcal{Q}\eta:\mathcal{Q}\rightarrow\mathcal{Q}\mathcal{H}$ is a morphism of monads. By Lemma 5.4, $(\mathcal{Q}\mathcal{H}(L),\rho_{L}\circ\mathcal{Q}\eta_{L})$ and $(L,\alpha\circ\mathcal{Q}\eta_{L})=(L,\alpha^{\mathcal{Q}})$ are $\mathcal{Q}$-algebras, and $\alpha:(\mathcal{Q}\mathcal{H}(L),\rho_{L}\circ\mathcal{Q}\eta_{L})\rightarrow(L,\alpha^{\mathcal{Q}})$ is a $\mathcal{Q}$-algebra homomorphism. By Proposition 2.4 (1), $(\mathcal{Q}\mathcal{H}(L),\iota_{\mathcal{H}(L)})$ is also a $\mathcal{Q}$-algebra. Since $\mathcal{Q}$ is a right $\rm{KZ}$-monad, Proposition 5.8 indicates that structure maps for $\rm{KZ}$-monads are unique, $\rho_{L}\circ\mathcal{Q}\eta_{L}=\iota_{\mathcal{H}(L)}$ holds; therefore, $\alpha:(\mathcal{Q}\mathcal{H}(L),\iota_{\mathcal{H}(L)})\rightarrow(L,\alpha^{\mathcal{Q}})$ is a $\mathcal{Q}$-algebra homomorphism.
  2. 2.
    Theorem 5.3 and Proposition 5.5 imply that $\theta_{\mathcal{H}}:\mathcal{H}\rightarrow\mathcal{Q}\mathcal{H}$ is a morphism of monads. By Lemma 5.4, $(\mathcal{Q}\mathcal{H}(L),\rho_{L}\circ\theta_{\mathcal{H}(L)})$ and $(L,\alpha\circ\theta_{\mathcal{H}(L)})=(L,\alpha^{\mathcal{H}})$ are $\mathcal{H}$-algebras, and $\alpha:(\mathcal{Q}\mathcal{H}(L),\rho_{L}\circ\theta_{\mathcal{H}(L)})\rightarrow(L,\alpha^{\mathcal{Q}})$ is an $\mathcal{H}$-algebra homomorphism.
  3. 3.
    Assume that $\beta:\mathcal{QH}(L)\rightarrow L$ is also a structure map of $L$. Then $\beta^{\mathcal{Q}}=\beta\circ\mathcal{Q}\eta_{L}$ and $\beta^{\mathcal{H}}=\beta\circ\theta_{\mathcal{H}(L)}$ are $\mathcal{Q}$-algebra and $\mathcal{H}$-algebra structure maps, respectively. Since both $\mathcal{Q}$ and $\mathcal{H}$ are $\rm{KZ}$-monads, it follows from Proposition 5.8 that each $\mathcal{Q}$-algebra ($\mathcal{H}$-algebra) has at most one structure map. Thus $\beta^{\mathcal{Q}}=\alpha^{\mathcal{Q}}$ and $\beta^{\mathcal{H}}=\alpha^{\mathcal{H}}$. By (1), $\alpha$ and $\beta$ are $\mathcal{Q}$-algebra homomorphisms from $(\mathcal{Q}\mathcal{H}(L),\iota_{\mathcal{H}(L)})$ to $(L,\alpha^{\mathcal{Q}})=(L,\beta^{\mathcal{Q}})$. Since $\alpha\circ\theta_{\mathcal{H}(L)}=\alpha^{\mathcal{H}}=\beta^{\mathcal{H}}=\beta\circ\theta_{\mathcal{H}(L)}$, Proposition 2.4 (3) implies $\alpha=\beta$.
Proposition 5.12.
If $L$ is a $\mathcal{QH}$-algebra, then $L$ is a frame.
Proof.
The fact given in Proposition 5.11 that $L$ is a $\mathcal{Q}$-algebra and an $\mathcal{H}$-algebra yields that $L$ is indeed a meet continuous complete lattice by Corollary 3.8 and Proposition 3.16. Thus the proof will be completed if we prove that $L$ is distributive.
$\mathbf{Claim~1}$: $\mathcal{QH}(L)$ is a distributive lattice.
Obviously, the union of arbitrary finite elements in $\mathcal{QH}(L)$ is still compact saturated. By Lemma 4.2, we know the complete lattice $\mathcal{H}(L)$ endowed with the Scott topology is coherent, that is, the arbitrary finite intersection of compact saturated subsets of $\mathcal{H}(L)$ is still in $\mathcal{QH}(L)$. Thus $\mathcal{QH}(L)$ is a lattice, in which the supremum and infimum of each finite subset are their union and intersection, respectively. This contributes to the distributivity of $\mathcal{QH}(L)$.
$\mathbf{Claim~2}$: The structure map $\alpha:\mathcal{QH}(L)\rightarrow L$ is a lattice homomorphism.
By Proposition 5.11, $\alpha$ is both a $\mathcal{Q}$-homomorphism and an $\mathcal{H}$-homomorphism. Thus $\alpha$ preserves finite infs by Lemma 3.15 and all sups by Lemma 3.7.
Now we consider the functions $\alpha\circ\varphi_{i}\circ(\theta_{\mathcal{H}}\circ\eta_{L})^{3}:L\times L\times L\rightarrow L,\ i=1,2$, where $(\theta_{\mathcal{H}}\circ\eta_{L})^{3}:L\times L\times L\rightarrow\mathcal{QH}(L)\times\mathcal{QH}(L)\times\mathcal{QH}(L)$ is defined as
$(\theta_{\mathcal{H}}\circ\eta_{L})^{3}(x,y,z)=({\uparrow}\{{\downarrow}x\},{\uparrow}\{{\downarrow}y\},{\uparrow}\{{\downarrow}z\})$, for any $x,y,z\in L$,
and $\varphi_{i}:(\mathcal{QH}(L))^{3}\rightarrow\mathcal{QH}(L),\ i=1,2$ are defined below: for each triple $(\mathcal{K}_{1},\mathcal{K}_{2},\mathcal{K}_{3})\in(\mathcal{QH}(L))^{3}$,
$\varphi_{1}(\mathcal{K}_{1},\mathcal{K}_{2},\mathcal{K}_{3})=\mathcal{K}_{1}\wedge(\mathcal{K}_{2}\vee\mathcal{K}_{3})$,
$\varphi_{2}(\mathcal{K}_{1},\mathcal{K}_{2},\mathcal{K}_{3})=(\mathcal{K}_{1}\wedge\mathcal{K}_{2})\vee(\mathcal{K}_{1}\wedge\mathcal{K}_{3})$.
One can calculate that $\alpha\circ\varphi_{1}\circ(\theta_{\mathcal{H}}\circ\eta_{L})^{3}(x,y,z)=x\wedge(y\vee z)$ and $\alpha\circ\varphi_{2}\circ(\theta_{\mathcal{H}}\circ\eta_{L})^{3}(x,y,z)=(x\wedge y)\vee(x\wedge z)$. It is obvious that $\varphi_{1}=\varphi_{2}$ by Claim 1. Thus $x\wedge(y\vee z)=(x\wedge y)\vee(x\wedge z)$, that is, $L$ satisfies the distributive law. In conclusion, $L$ is a frame.
Proposition 5.13.
Every $\mathcal{QH}$-algebra homomorphism is a frame homomorphism.
Proof.
Let $f:(L,\alpha)\rightarrow(M,\beta)$ be a $\mathcal{QH}$-algebra homomorphism. It was shown in Theorem 5.3 that $\mathcal{Q}\eta:\mathcal{Q}\rightarrow\mathcal{QH}$ and $\theta_{\mathcal{H}}:\mathcal{H}\rightarrow\mathcal{QH}$ are morphisms of monads; hence, by Lemma 5.4, we know $f:(L,\alpha\circ\mathcal{Q}\eta_{L})\rightarrow(M,\beta\circ\mathcal{Q}\eta_{M})$ is a $\mathcal{Q}$-algebra homomorphism and $f:(L,\alpha\circ\theta_{\mathcal{H}(L)})\rightarrow(M,\beta\circ\theta_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{H}(M)}})$ is an $\mathcal{H}$-algebra homomorphism. It follows that $f$ preserves finite infs and arbitrary sups, that is to say, $f$ is a frame homomorphism.
Let $\mathbf{Frm}$ be the category of all frames and all frame homomorphisms.
Theorem 5.14.
The Eilenberg-Moore category of $\mathcal{QH}$ over $\mathbf{WF}$ is a subcategory of the category $\mathbf{Frm}$.
Remark 5.15.
See the deflationary semilattice in $\mathbf{WF}$ quoted in Example 3.17 $\mathcal{L}$ again, it has been proven not to be a $\mathcal{Q}$-algebra. Thus evidenced by Proposition 5.11, $\mathcal{L}$ is not a $\mathcal{QH}$-algebra, which indicates that $\mathcal{QH}$ cannot give the free deflationary semilattices for $\mathbf{WF}$.
Remark 5.16.
The general observation that the algebras of the double power construction are frames is due to S. Vickers [24] or perhaps P. Taylor [23], although Vickers works with locales and Taylor works in an abstract category of locally compact sober spaces. In an abstract category of locally compact sober spaces, Taylor proved that when the power constructions commute, the composite monad is equivalent to the monad defined as double exponentiation by the Sierpinski space, and characterized its algebras as frames. Furthermore, the characterizations work in more general categories when the necessary exponentials exist. We know every core-compact well-filtered dcpo is locally compact sober [17] and by Corollary 4.9, for each locally compact sober dcpo, the Hoare and the Smyth power constructions commute with each other. Thus, our composite monad $\mathcal{Q}\mathcal{H}$ is also equivalent to the monad defined as double exponentiation by the Sierpinski space in the core-compact case. Then a natural open question is whether this equivalence extends to a more general category of well-filtered dcpos satisfying conditions (i) and (ii) of Theorem 4.9. We leave this for future investigation.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.